### v3

# DAS_CE_Supp

Abstract
By deploying a large number of antennas with sub-half-wavelength spacing in a compact space, dense array systems (DASs) can fully unleash the multiplexing and diversity gains of limited apertures. To acquire these gains, accurate channel state information acquisition is necessary but challenging due to the large antenna numbers. To overcome this obstacle, this paper reveals that designing the observation matrix to exploit the high spatial correlation of DAS channels is crucial for realizing near-optimal Bayesian channel estimation. Specifically, we prove that the observation matrix design for channel estimation is equivalent to a time-domain duality of point-to-point multiple-input multiple-output precoding, except for the change in the total power constraint on the precoding matrix to the pilot-wise discrete power constraint on the observation matrix. Inspired by Bayesian regression, a novel ice-filling algorithm is proposed to design amplitude-and-phase controllable observation matrices, and a majorization-minimization algorithm is proposed to address the phase-only controllable case. Particularly, we prove that the ice-filling algorithm can be interpreted as a “quantized” water-filling algorithm, wherein the latter’s continuous power-allocation process is converted into the former’s discrete pilot-assignment process. To support the near-optimality of the proposed designs, we provide comprehensive analyses on the achievable mean square errors and their asymptotic expressions. Finally, numerical results confirm that our proposed designs achieve the near-optimal channel estimation performance and outperform existing approaches significantly.
Index Terms:
Estimation theory, mutual-information maximization, dense array systems (DAS), Bayesian regression.
I Introduction
From 3G to 5G, antenna array systems play an irreplaceable role in wireless communications [1, 2]. By strategically configuring multiple antennas, an antenna array can constructively manipulate the radiated signals, allowing for enhanced wireless coverage, improved spectral efficiency, and reduced power consumption [3, 4, 5]. The performance of arrays continuously improves with the number of antennas [6, 7]. However, in practical scenarios, the space available for antenna deployment is usually limited, which restricts the performance gains endowed by antenna arrays [8]. To achieve a better performance with a spatially-limited aperture, dense array systems have attracted extensive attentions in recent years [9].
Generally, DASs represent a series of array technologies that massive sub-wavelength antennas are densely arranged within a compact space. Unlike the conventional arrays with half-wavelength antenna spacing $\lambda/2$, the antenna spacing of DASs is much smaller, such as $\lambda/6$ [10], $\lambda/8$ [11], $\lambda/10$ [12], or even $\lambda/23$ [13]. With an increased number of sub-channels, DASs promise to achieve high array gains and fully exploit the multiplexing and diversity gains of limited apertures [9]. Besides, DASs can also reduce the grating lobes and provide high performance for large values of oblique angles of incidence [14]. The typical DAS realizations include holographic multiple-input multiple-output (MIMO) [15], reconfigurable intelligent surfaces [16], fluid antenna systems (FASs) [17, 18], graphene-based nano-antenna arrays [19, 20], radio-frequency lens [21], and so on. For example, in [10], massive sub-wavelength patches of $\lambda/6$-spacing are densely printed on a holographic MIMO surface to generate multiple beams flexibly. In [22], meta-elements sized of $\lambda/2.5\times\lambda/17.5$ are seamlessly integrated onto a feeding microstrip to achieve reconfigurable holographic surface (RHS) beam steering. In [13], a RIS composed of antennas sized of $\lambda/23\times\lambda/23$ is designed and fabricated for terahertz communications.
The performance gains of antenna arrays are realized via the constructive beamformers enabled by their phase-shifters (PSs) and radio frequency (RF) chains. To realize effective beamforming, the acquisition of accurate channel state information (CSI) is essential, particularly in situations where the number of RF chains is smaller than the number of massive antennas. Up to now, many estimators have been proposed to acquire the CSI of large antenna arrays, and most of them can be adopted in DASs [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. Their fundamental principle involves receiving pilot signals using observation matrices and recovering the channel by advanced estimation algorithms. For example, when the available pilot length is larger than the antenna number, the classical least square (LS) estimator is usually adopted. Leveraging the property of channel sparsity, compressed sensing (CS)-based channel estimators are widely studied to improve the estimation accuracy and reduce the pilot overhead [23, 24], such as the orthogonal matching pursuit (OMP)-based estimator [25, 26], the message passing (MP)-based estimator [27, 28, 29], and the gridless sparse signal reconstructor [30]. By training neural networks with a large amounts of channel data, the deep learning approaches are also utilized to realize both data-driven and model-driven channel estimators [31, 32, 33, 37]. Besides, beam alignment techniques [34, 35, 36], including beam sweeping and hierarchical beam training, have also been widely explored to acquire the implicit CSI with low pilot overhead.
Although most of existing channel estimators can be adopted in DASs, they fail to fully exploit the high spatial correlation of DAS channels, thereby leaving a remarkable performance gap from the optimal estimator. Specifically, this spatial correlation is attributed to the fact that, the extremely-dense deployment of DAS antennas significantly increases the similarity of radio waves impinging on antenna ports and aggravates the mutual-coupling effect between adjacent antenna circuits. This fact makes the covariance matrices of DAS channels no longer diagonal but highly structured. For diagonal covariance matrices, the observation matrices for receiving pilot signals are either generated randomly or set as predefined codebooks, such as the Discrete Fourier Transform (DFT) matrix [23, 24, 25, 30, 32, 27, 26, 28, 31, 37, 34, 35, 29, 33, 36], which are sufficient to achieve the optimality of channel estimation [38, 39]. In contrast, since the correlation matrix of DAS channels are full of structural features, it is believed that their observation matrices can be tightly aligned with these features in pilot transmission. This structured pilot-sensing process has a high potential to remarkably boost the channel estimation accuracy in DASs [40, 41, 42]. For example, in [18], the structured channel covariance is utilized to determine the positions of fluid antennas during the pilot reception. Due to the strong channel correlation of FAS channels, the estimator in [18] can achieve higher accuracy than CS methods. Authors in [43] further extend the method in [18] to sparse channel scenarios. In [44], aiming at efficient CSI acquisition in holographic MIMO systems, the electromagnetic channel covariance is modeled and utilized to enable a Bayesian channel estimator. However, these existing estimators often rely on specific array structures or parametric channel models, which can neither be used to design the general observation matrix in DASs nor guarantee the performance to be optimal.
To fill in this blank, our work represents the first attempt on designing and analyzing the near-optimal observation matrices in the DAS channel estimation. The classical Bayesian regression method is adopted as the channel estimator, i.e., the minimum-mean-square-error (MMSE) method, while our key contributions lie in the designed observation matrices to enhance Bayesian regression schemes, as summarized below.
  • Mutual-information-maximization (MIM) based framework: From the perspective of MIM, we propose to design the DAS’s observation matrix by minimizing the information uncertainty between the received pilots and wireless channels. The formulated MIM problem for observation matrix design is shown to be a time-domain duality of point-to-point MIMO precoding design. In particular, the multi-RF-chain resources for MIMO precoding is transformed to the multi-timeslot-pilot resources for channel estimation, and the total power constraint on the precoding matrix is switched to the pilot-wise discrete power constraint on the observation matrix. Motivated by this finding, we show that the ideal (but unachievable in practice) observation matrix could be constructed by the water-filling principle, which involves the use of the eigenvectors of channel covariance matrix as well as optimal power allocation.
  • Ice-filling enabled observation matrix design: For practical DAS channel estimations, the continuous power-allocation by water-filling is not applicable as the pilot length is discrete and each pilot transmission has independent power constraint. To deal with this issue, we propose an ice-filling algorithm to design the amplitude-and-phase controllable observation matrices. Inspired by Bayesian regression, this algorithm sequentially generates the optimal blocks of the observation matrix by maximizing the mutual information (MI) increment between two adjacent pilot transmissions. An important insight is that the ice-filling algorithm converts the power-allocation-process of the ideal water-filling algorithm into an eigenvector-assignment-process. We rigorously prove that, the continuous powers allocated by water-filling are quantized as the number of times that each eigenvector of the channel covariance is assigned for pilot transmission by ice-filling, with a quantization error smaller than one. The proven quantization nature thereby ensures the near-optimality of the proposed ice-filling algorithm.
  • Majorization minimization (MM) enabled observation matrix design: We then apply our framework for DAS channel estimation to the situation when only the phases of the observation matrix’s weights are controllable while their amplitudes are fixed, making both the water-filling and ice-filling algorithms invalid. To address this challenge, we propose a MM algorithm for designing the observation matrix. Its novelty lies in replacing the primal non-convex MIM problem with a series of tractable approximate subproblems having analytical solutions, which are solved in an alternating optimization manner. Numerical simulations demonstrate that its channel estimation performance is comparable to the ice-filling algorithm.
  • Performance analysis: Comprehensive analyses on the achievable mean square errors (MSEs) are provided to validate the effectiveness of the proposed designs. We prove that, compared to the random observation matrix design, the ice-filling algorithm can significantly improve the estimation accuracy by the order of the ratio between the number of antennas and the rank of channel covariance. Furthermore, based on the derived quantization nature of ice-filling, the MSE gap between water-filling and ice-filling is proved to decay quadratically with the pilot length, demonstrating the near-optimality of the ice-filling algorithm. In addition, the close-form MSEs under imperfect channel covariance are also derived. The superiority of the proposed algorithms is further validated using extensive numerical results.
The remainder of the paper is organized as follows. The system model is presented in Section II. The proposed MIM-based framework for observation matrix design is given in Section III. The ice-filling and MM algorithms are elaborated in Section IV. Then, the MSEs are analyzed in Section V. Simulation results are carried out in Section VI, and finally the conclusions are drawn in Section VII.
Notation: Lower-case and upper-case boldface letters represent vectors and matrices, respectively. ${[\cdot]^{-1}}$, ${[\cdot]^{{\dagger}}}$, ${[\cdot]^{*}}$, ${[\cdot]^{\rm T}}$, and ${[\cdot]^{\rm H}}$ denote the inverse, pseudo-inverse, conjugate, transpose, and conjugate-transpose operations, respectively; $\|\cdot\|_{2}$ denotes the $l_{2}$-norm of the argument; $\|\cdot\|_{F}$ denotes the Frobenius norm of the argument; ||$\cdot$|| ⋅ | denotes the element-wise amplitude of its argument; $\mathsf{Tr}(\cdot)$ denotes the trace operator; ${\mathsf{E}}\left(\cdot\right)$ is the expectation operator; ${}\Re\{\cdot\}$ denotes the real part of the argument; $\lambda_{\max}(\cdot)$ denotes the largest eigenvalue of its argument; $\mathcal{C}\mathcal{N}\!\left({\bm{\mu}},{\bf\Sigma}\right)$ denotes the complex Gaussian distribution with mean ${\bm{\mu}}$ and covariance ${\bf\Sigma}$; ${\cal U}\left(a,b\right)$ denotes the uniform distribution between $a$ and $b$; $\mathbf{I}_{L}$ is an $L\times L$ identity matrix; $\mathbf{1}_{L}$ is an $L$-dimensional all-one vector; and $\mathbf{0}_{L}$ is an all-zero vector or matrix with dimension $L$.
II System Model
Figure 1: An illustration of uplink channel estimation for a DAS.
As illustrated in Fig. 1, we consider the narrowband channel estimation of an uplink single-input multiple-output (SIMO) system. The base station (BS) employs an $M$-antenna DAS, which comprises an analog combining structure supported by one RF chain, to receive the pilots from a single-antenna user.
Let ${\bf h}\in{\mathbb{C}}^{M\times 1}$ be the channel vector and $Q$ the number of transmit pilots within a coherence-time frame. The received signal ${y}_{q}\in{\mathbb{C}}$ at the BS in timeslot $q$ is modeled as
${y}_{q}={\bf w}^{H}_{q}{\bf h}s_{q}+{\bf w}^{H}_{q}{\bf z}_{q},$ (1)
where ${\bf w}_{q}\in{\mathbb{C}}^{M\times 1}$ is the observation vector at the BS, $s_{q}$ the pilot transmitted by the user, and ${\bf z}_{q}\sim\mathcal{C}\mathcal{N}\!\left({\bf 0}_{M},\sigma^{2}{\bf I}_{M}\right)$ the additive white Gaussian noise. As the observation vector ${\bf w}_{q}$ affects both the desired signal ${\bf h}s_{q}$ and the noise ${\bf z}_{q}$, its power has no impact on the estimation accuracy. Therefore, without any loss of generality, we can normalize ${\bf w}_{q}$ to $\|{\bf w}_{q}\|^{2}_{2}=1$, which reshapes the noise distribution to ${\bf w}_{q}^{H}{\bf z}_{q}\sim\mathcal{C}\mathcal{N}\!\left(0,\sigma^{2}\right)$. Two practical implementations of the observation vector ${\bf w}_{q}$ will be discussed in our research: the amplitude-and-phase controllable combiner and the phase-only controllable combiner. The former connects each antenna to the RF chain using one PS and one low noise amplifier (LNA), enabling adjustment of both the amplitude and phase of the coefficients of ${\bf w}_{q}$. The latter deploys a single PS between each antenna and the RF chain, allowing only phase adjustment of the coefficients of ${\bf w}_{q}$ and enforcing a constant modulus constraint on the observation vector, i.e., $|w_{q,m}|=\frac{1}{\sqrt{M}}$ for $m\in\{1,2,\cdots,M\}$ [23].
Considering the total $Q$-timeslot pilot transmission, the received signal vector ${\bf y}:=\left[{y}_{1},\cdots,{y}_{Q}\right]^{\rm T}$ is expressed as
${\bf y}={\bf W}^{H}{\bf h}+{\bf z},$ (2)
where $s_{q}$ is assumed to be 1 for all $q\in\{1,\cdots,Q\}$, ${\bf{W}}=\left[{\bf{w}}_{1},\cdots,{\bf{w}}_{Q}\right]$ stands for the observation matrix, and ${\bf z}:=\left[{\bf w}^{H}_{1}{\bf z}_{1},\cdots,{\bf w}^{H}_{Q}{\bf z}_{Q}\right]^{\rm T}$. As the power of ${\mathbf{w}}_{q}$ is normalized to 1, we have ${\bf z}\sim\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$. The objective of channel estimation is to recover the channel $\bf h$ from the received pilot $\bf y$. As a DAS possesses highly correlated channels across different antenna ports, we adopt Bayesian regression to recover ${\mathbf{h}}$. Suppose the channel is generated from a Gaussian process $\mathcal{CN}({\mathbf{0}}_{M},\bm{\Sigma}_{\bf h})$, where the covariance matrix $\bm{\Sigma}_{\bf h}\in\mathbb{C}^{M\times M}$, also called prior kernel, characterizes our prior knowledge to the channel correlation. The kernel reflects the long-term statistical information of the covariance of channels, i.e., ${\bf\Sigma}={\mathsf{E}}({\bf h}{\bf h}^{H})$, which can keep unchanged for a long time. To obtain the prior kernel in practical MIMO systems, an ideal approach is to leverage the channel realizations generated by some standard models or real-world datasets. In addition, some online second-moment learning methods proposed in [45, 46] can also be used to obtain the prior kernel efficiently. Given the prior kernel, the joint probability distribution of $[{\mathbf{h}}^{T},{\mathbf{y}}^{T}]^{T}$ satisfies
[…]
By invoking the property of the linear transform of Gaussian random vectors, the posterior mean and posterior covariance of ${\mathbf{h}}$ given observation ${\mathbf{y}}$ can be calculated as:
${{\bm{\mu}}}_{{\mathbf{h}}|{\mathbf{y}}}$ $={{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}}{\left({\bf W}^{H}{\bf{\Sigma}}_{\bf{h}}{\bf W}+\sigma^{2}{\bf I}_{Q}\right)^{-1}}{{\bf{y}}},$ (4)
${{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}$ $={{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}}{\left({\bf W}^{H}{\bf{\Sigma}}_{\bf{h}}{\bf W}+\sigma^{2}{\bf I}_{Q}\right)^{-1}}{{{{\bf{W}}^{H}}}}{{\bf{\Sigma}}_{\bf{h}}}.$ (5)
The optimal Bayesian estimation to the channel vector is exactly the posterior mean, i.e., $\hat{{\mathbf{h}}}={{\bm{\mu}}}_{{\mathbf{h}}|{\mathbf{y}}}$, which is also known as the MMSE estimator. Notably, the posterior covariance, ${\bf\Sigma_{{\mathbf{h}}|{\mathbf{y}}}}$, also called posterior kernel, characterizes the channel estimation error and is a function of the prior kernel ${{\bf{\Sigma}}_{\bf{h}}}$ and the observation matrix ${\mathbf{W}}$. Due to the extremely high spatial correlation exhibited by DAS channels, the prior kernel ${{\bf{\Sigma}}_{\bf{h}}}$ would remarkably deviate from the identity matrix. This deviation indicates that aligning the observation matrix ${\mathbf{W}}$ with the kernel’s subspace could be greatly helpful in reducing the estimation error. Motivated by this fact, this paper concentrates on the design of observation matrix ${\mathbf{W}}$, by considering both the amplitude-and-phase controllable and phase-only controllable analog combiners, to boost the channel estimation performance.
III Mutual Information Maximization for Channel Estimation in DASs
In this section, we first present a MIM based framework for designing observation matrices. Then, a water-filling-inspired observation-matrix-design is investigated as an ideal case of the proposed framework.
III-A MIM based framework for observation matrix design
We adopt the MI between the received signal ${\mathbf{y}}$ and the channel ${\mathbf{h}}$ as the metric to evaluate the quality of the designed observation matrix. There are two reasons for using the MI as a metric. Firstly, MI tells the amount of information about the unknown channel ${\mathbf{h}}$ we can gain from the received signal ${\mathbf{y}}$. The second reason is that, according to the information-theoretic properties of Bayesian statistics from [47], the maximization of MI is asymptotically the minimization of Cramer Rao Bound (CRB).
Given the distributions ${\mathbf{h}}\sim\mathcal{CN}({\mathbf{0}}_{M},\bm{\Sigma}_{\bf h})$ and ${\bf z}\sim\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$, the MI is formulated as
$\max_{{\mathbf{W}}\in\mathcal{W}}I({\mathbf{y}};{\mathbf{h}})$ $=\log\det\left({\mathbf{I}}_{Q}+\frac{1}{\sigma^{2}}{\mathbf{W}}^{H}{\bf\Sigma_{h}}{\mathbf{W}}\right),$ (6)
where $I(\cdot;\cdot)$ denotes the MI, $\det(\cdot)$ is the determinant of its argument, and $\mathcal{W}$ represents the feasible set of ${\mathbf{W}}$ for both amplitude-and-phase controllable and phase-only controllable analog combiners. It is notable that the observation matrix design problem in (6) resembles the well-know point-to-point MIMO precoding problem. Comprehensive comparisons between these two problems are provided below.
  • Category of unknown data: Both MIMO precoding and observation matrix design in (6) aim to reduce the uncertainty between the received data, ${\mathbf{y}}$, and the unknown data. The former regards the transmitted symbol vector, denoted by ${\mathbf{x}}$, as the unknown data. To facilitate MIMO precoding, multiple RF chains are employed to support an equal number of data streams. On the other hand, the unknown data to the latter is the channel, ${\mathbf{h}}$. Our considered uplink channel estimation requires transmitting pilot signals over multiple timeslots to recover ${\mathbf{h}}$. In comparison, one can discover that the multiple RF chains in MIMO precoding are converted to the multi-timeslot pilot resources in channel estimation. Thereby, the design of the observation matrix in (6) can be viewed as a time-domain duality of point-to-point MIMO precoding.
  • Constraint on the observation matrix: The mathematical difference between MIMO precoding and observation matrix design for uplink channel estimation is attributed to the constraints of ${\mathbf{W}}$. For MIMO precoding, each column of ${\mathbf{W}}$ refers to a precoding vector associated with one dedicated RF chain. The optimization of precoding vectors, leveraging multiple RF chains, typically enforces the total-power-constraint on ${\mathbf{W}}$, i.e., $\|{\mathbf{W}}\|_{F}^{2}=Q$. In terms of the observation matrix design, each column of ${\mathbf{W}}$ signifies an observation vector used for receiving one pilot signal. As discussed in Section II, since each observation vector ${\mathbf{w}}_{q}$ amplifies both the desired signal and the noise, it is subject to the pilot-wise power constraint, i.e., $\|{\mathbf{w}}_{q}\|_{2}^{2}=1$ or $|w_{q,m}|=\frac{1}{\sqrt{M}}$. The distinction in the observation matrix constraint differentiates the MIM in (6) from the classical MIMO precoding.
III-B Water-filling Inspired Ideal Observation Matrix Design
Prior to addressing problem (6), we would like to consider an ideal (but practically unachievable) situation to establish an upper bound of (6). Specifically, we temporarily make the ideal assumption that the noise vector ${\mathbf{z}}$ in (2) is independent of the observation matrix ${\mathbf{W}}$ and its distribution is always $\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$ regardless of the power of ${\mathbf{w}}_{q}$. In this context, the pilot-wise power constraint $\|{\mathbf{w}}_{q}\|^{2}_{2}=1$ can be relaxed to the total power constraint $\|{\mathbf{W}}\|_{F}^{2}=\sum_{q=1}^{Q}\|{\mathbf{w}}_{q}\|^{2}_{2}=Q$, and problem (6) becomes identical to the point-to-point MIMO precoding problem, which can be solved by water-filling [48].
Let the eigenvalue decomposition (EVD) of ${\bf\Sigma_{h}}$ be ${\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}$, where $K$ is the rank of $\bf\Sigma_{h}$, ${\mathbf{U}}_{K}=[{\mathbf{u}}_{1},{\mathbf{u}}_{2},\cdots,{\mathbf{u}}_{K}]$, and ${\bf\Lambda}_{K}={\rm diag}\{{\lambda}_{1},{\lambda}_{2},\cdots,{\lambda}_{K}\}$ with ${\lambda}_{1}\geq\cdots\geq\lambda_{K}>0$. The ideal observation matrix ${\mathbf{W}}^{\rm Ideal}$ is given as ${\mathbf{W}}^{\rm Ideal}={\mathbf{U}}_{K}{\mathbf{P}}$. The power allocation matrix ${\mathbf{P}}\in\mathbb{R}^{K\times Q}$ is expressed as ${\mathbf{P}}=[{\mathbf{P}}_{K},{\mathbf{0}}_{K\times(Q-K)}]$ with ${\mathbf{P}}_{K}={\rm diag}\{\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\}$. The power $p_{k}$ allocated to each eigenvector is calculated by the water-filling principle:
$p_{k}=(\beta-{\sigma^{2}}/{\lambda_{k}})^{+},\forall k\in\{1,2,\cdots,K\},$ (7)
where $(x)^{+}\overset{\Delta}{=}\max\{x,0\}$, and the water-level $\beta$ is properly selected to meet the power constraint: $\sum_{k=1}^{K}p_{k}=Q$.
IV Practical Observation Matrix Design
This section elaborates on the observation matrix design in practical scenarios. Firstly, a greedy-method based principle is leveraged to solve the MIM problem in (6). Building on this principle, we propose an ice-filling algorithm for designing the amplitude-and-phase controllable observation matrices. Additionally, we conduct a comprehensive analysis on the relationship between ice-filling and water-filling. Subsequently, a MM-based algorithm is proposed for the design of phase-only controllable observation matrices. Last, the choice of prior kernel is also discussed.
IV-A Observation Matrix Design Using Greedy Method
Figure 2: Framework of the proposed observation matrix design.
In practical uplink channel estimation where the noise vector ${\mathbf{z}}$ is amplified by the observation vector ${\mathbf{w}}_{q}$, the pilot-wise power constraint should be considered. In this case, it is intractable to find the optimal solution to (6). To overcome this challenge, we harness a greedy method to obtain the observation vectors in a pilot-by-pilot manner, as shown in Fig. 2. Define ${\bf{W}}_{t}=[{\bf{w}}_{1},{\bf{w}}_{2},\cdots,{\bf{w}}_{t}]$ as the observation matrix for timeslots $1\sim t$, where $t<Q$, and denote ${\mathbf{y}}_{t}={\mathbf{W}}_{t}^{H}{\mathbf{h}}+{\mathbf{z}}_{t}$ as the corresponding received signal. Given the current observation matrix ${\bf{W}}_{t}$, our greedy method aims to determine the next observation vector, ${\mathbf{w}}_{t+1}$, by maximizing the MI increment from timeslot $t$ to $t+1$, i.e., $\max_{{\mathbf{w}}_{t+1}}I({\mathbf{y}}_{t+1};{\mathbf{h}})-I({\mathbf{y}}_{t};{\mathbf{h}})$. As proved in [49], the MI increment is expressed as
$I({\mathbf{y}}_{t+1};{\mathbf{h}})-I({\mathbf{y}}_{t};{\mathbf{h}})=\log_{2}\left(1+\frac{1}{\sigma^{2}}{\mathbf{w}}_{t+1}^{H}{\bm{\Sigma}}_{t}{\mathbf{w}}_{t+1}\right),$ (8)
where ${{\bf{\Sigma}}_{t}}$ represents the posterior kernel of ${\mathbf{h}}$ given the observation ${\mathbf{y}}_{t}$. Using the block matrix inverse, we can explicitly write ${\bf{\Sigma}}_{t}$ as a function of ${\bf{\Sigma}}_{t-1}$ and ${\mathbf{w}}_{t}$ (Detailed proof can be found in Appendix A):
${\bf{\Sigma}}_{t}={\bf{\Sigma}}_{t-1}-\frac{{\bf{\Sigma}}_{t-1}{\bf{w}}_{t}{\bf{w}}_{t}^{H}{\bf{\Sigma}}_{t-1}}{{\bf{w}}_{t}^{H}{\bf{\Sigma}}_{t-1}{\bf{w}}_{t}+\sigma^{2}}.$ (9)
Accordingly, the optimal ${\bf{w}}_{t+1}$ can be attained by solving the quadratic problem:
${\bf{w}}_{t+1}=\mathop{\arg\!\max}\limits_{{\bf{w}}\in\mathcal{W}}{\bf{w}}^{H}{\bf{\Sigma}}_{t}{\bf{w}}.$ (10)
As a result, equations (8) to (10) inspire us to propose the framework in Fig. 2 for designing observation matrices. It begins by setting ${\bm{\Sigma}}_{0}$ as the prior kernel, ${\bm{\Sigma}}_{\bf h}$. Then, the first observation vector ${\mathbf{w}}_{1}$ is obtained by solving problem (10). In subsequent timeslots $t\in\{1,2,\cdots,Q-1\}$, the posterior kernels ${\bm{\Sigma}}_{t}$ are updated from ${\bm{\Sigma}}_{t-1}$ and ${\bf w}_{t}$ using (9), while the observation vectors ${\bf w}_{t+1}$ are updated from ${\bm{\Sigma}}_{t}$ by solving problem (10). These two updating steps are repeated until all $Q$ observation vectors are generated for performing the ensuing Bayesian regression.
In the sequel, we will elaborate on solving problem (10) under the amplitude-and-phase controllable and phase-only controllable analog combiners, respectively.
IV-B Ice-filling for Amplitude-and-phase Controllable Observation Matrix
Algorithm 1 Ice-Filling-Based Observation Matrix Design
0:  Number of pilots $Q$, kernel ${\bm{\Sigma}}_{\bf h}$.
0:  Designed observation matrix ${\bf W}$.
1:  Find the eigenvectors $[{\mathbf{u}}_{1},{\mathbf{u}}_{2},\cdots,{\mathbf{u}}_{K}]$ and the corresponding eigenvalues $[\lambda_{1},\lambda_{2},\cdots,\lambda_{K}]$ of ${\bf\Sigma}_{0}={\bf\Sigma_{h}}$
2:  Initialize $[\lambda_{1}^{0},\lambda_{2}^{0},\cdots,\lambda_{K}^{0}]=[\lambda_{1},\lambda_{2},\cdots,\lambda_{K}]$
3:  for $t=0,\cdots,Q-1$ do
4:     $k_{t}=\arg\max_{k\in\{1,2,\cdots,K\}}\{\lambda_{k}^{t}\}$
5:     Eigenvector-assignment: ${\mathbf{w}}_{t+1}={\mathbf{u}}_{k_{t}}$
6:     Eigenvalue-update: $\lambda_{k_{t}}^{t+1}=\frac{\lambda_{k_{t}}^{t}\sigma^{2}}{\lambda_{k_{t}}^{t}+\sigma^{2}}$
7:     Eigenvalue-preserve: $\lambda_{k}^{t+1}=\lambda_{k}^{t}$ for $k\neq k_{t}$
8:  end for
9:  Construct observation matrix: ${\bf W}=[{\mathbf{w}}_{1},{\mathbf{w}}_{2},\cdots,{\mathbf{w}}_{Q}]$
10:  return  Designed observation matrix ${\bf W}$
For amplitude-and-phase controllable observation matrices, problem (10) is the well-known Rayleigh quotient problem:
${\bf{w}}_{t+1}=\mathop{\arg\!\max}\limits_{\|{\bf w}\|_{2}^{2}=1}{\bf{w}}^{H}{\bf{\Sigma}}_{t}{\bf{w}}.$ (11)
The principal eigenvector of ${\bf{\Sigma}}_{t}$ yields the optimal solution to ${\mathbf{w}}_{t+1}$, i.e., ${\bf{\Sigma}}_{t}{\bf{w}}_{t+1}=\lambda_{\max}({\bf{\Sigma}}_{t}){\bf{w}}_{t+1}$. By leveraging the updating strategy of ${\bm{\Sigma}}_{t}$ described in (9) and the property of eigenvector, we can prove Theorem 1, which lays the foundation for our ice-filling algorithm.
Theorem 1
If ${\mathbf{w}}_{t}$ is the principal eigenvector of ${\bm{\Sigma}}_{t-1}$, then ${\bm{\Sigma}}_{t}$ can be rewritten as
${\bf{\Sigma}}_{t}={\bf\Sigma}_{t-1}-\frac{\lambda_{\max}^{2}({\bf\Sigma}_{t-1})}{\lambda_{\max}({\bf\Sigma}_{t-1})+\sigma^{2}}{\mathbf{w}}_{t}{\mathbf{w}}_{t}^{H}.$ (12)
Proof:
Applying the property of the principal eigenvector ${\bf\Sigma}_{t-1}{\mathbf{w}}_{t}=\lambda_{\max}({\bm{\Sigma}}_{t-1}){\mathbf{w}}_{t}$, we get ${\bf{w}}_{t}^{H}{\bf{\Sigma}}_{t-1}{\bf{w}}_{t}=\lambda_{\max}({\bf\Sigma}_{t-1})$ and ${\bf{\Sigma}}_{t-1}{\bf{w}}_{t}{\bf{w}}_{t}^{H}{\bf{\Sigma}}_{t-1}=\lambda_{\max}^{2}({\bf\Sigma}_{t-1}){\mathbf{w}}_{t}{\mathbf{w}}^{H}_{t}$, which together with (9) give rise to (12). ∎
Remark 1
Given that ${\mathbf{w}}_{t}$ is the principal eigenvector of ${\bm{\Sigma}}_{t-1}$, Theorem 1 implies that the posterior kernels, ${\bm{\Sigma}}_{t}$ and ${\bm{\Sigma}}_{t-1}$, share the identical eigenspace. This fact immediately leads to the result that the eigenvectors, ${\mathbf{U}}_{K}=[{\mathbf{u}}_{1},{\mathbf{u}}_{2},\cdots,{\mathbf{u}}_{K}]$, of the prior kernel, ${\bm{\Sigma}}_{0}={\bm{\Sigma}}_{{\mathbf{h}}}$, are inherited by all subsequent posterior kernels, ${\bm{\Sigma}}_{1},{\bm{\Sigma}}_{2},\cdots,$ and ${\bm{\Sigma}}_{Q-1}$. Thereby, we can draw the conclusion that all observation vectors, $\{{\mathbf{w}}_{1}\}_{t=1}^{Q}$, are selected from the eigenvectors of ${\bm{\Sigma}}_{{\mathbf{h}}}$.
Remark 2
Consider the eigenvalues of ${\bf\Sigma}_{t-1}$ and ${\bf\Sigma}_{t}$. We denote the EVD of ${\bf\Sigma}_{t-1}$ as ${\bf\Sigma}_{t-1}={\mathbf{U}}_{K}{\rm diag}\{\lambda_{1}^{t-1},\lambda_{2}^{t-1},\cdots,\lambda_{K}^{t-1}\}{\mathbf{U}}_{K}^{H}$, where $\{\lambda_{k}^{t-1}\}_{k=1}^{K}$ refers to the eigenvalues of $\{{\mathbf{u}}_{k}\}_{k=1}^{K}$, respectively. We use $k_{t-1}$ to index the principal eigenvalue of ${\bf\Sigma}_{t-1}$. Accordingly, the equations ${\mathbf{w}}_{t}={\mathbf{u}}_{k_{t-1}}$ and $\lambda_{\max}({\bf\Sigma}_{t-1})=\lambda_{k_{t-1}}^{t-1}$ hold, and Theorem 1 can be rewritten as:
${\bf{\Sigma}}_{t}$ $={\mathbf{U}}_{K}{\rm diag}\{\lambda_{1}^{t-1},\cdots,\lambda_{k_{t-1}}^{t-1},\cdots,\lambda_{K}^{t-1}\}{\mathbf{U}}_{K}^{H}$
$\quad\quad\quad\quad\quad\quad-\frac{(\lambda_{k_{t-1}}^{t-1})^{2}}{\lambda_{k_{t-1}}^{t-1}+\sigma^{2}}{\mathbf{u}}_{k_{t-1}}{\mathbf{u}}_{k_{t-1}}^{H}$ (13)
$={\mathbf{U}}_{K}{\rm diag}\left\{\lambda_{1}^{t-1},\cdots,\frac{\lambda_{k_{t-1}}^{t-1}\sigma^{2}}{\lambda_{k_{t-1}}^{t-1}+\sigma^{2}},\cdots,\lambda_{K}^{t-1}\right\}{\mathbf{U}}_{K}^{H}.$
Equation (2) reveals that only the principal eigenvalue, $\lambda_{k_{t-1}}^{t-1}$, of ${\bm{\Sigma}}_{t-1}$ is squeezed to an eigenvalue of ${\bm{\Sigma}}_{t}$ given by $\frac{\lambda_{k_{t-1}}^{t-1}\sigma^{2}}{\lambda_{k_{t-1}}^{t-1}+\sigma^{2}}$, while the remaining eigenvalues, $\lambda_{k}^{t-1},\forall k\in\{1,2,\cdots,K\}/\{k_{t-1}\}$, keep unchanged.
Motivated by Remark 1 and Remark 2, our ice-filling algorithm is intrinsically an eigenvector-assignment-process, as presented in Algorithm 1. To elaborate, the eigenvalues $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$ of the posterior kernel ${\bm{\Sigma}}_{t}$ are initialized as the eigenvalues of the prior kernel ${\bm{\Sigma}}_{\bf h}$ at $t=0$ in Step 2. In each timeslot, we find the largest eigenvalue from $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$ and index it by $k_{t}$ in Step 4. Then, the corresponding eigenvector ${\mathbf{u}}_{k_{t}}$ of ${\bm{\Sigma}_{h}}$, which is equivalent to the principal eigenvector of ${\bm{\Sigma}}_{t}$, is assigned to ${\mathbf{w}}_{t+1}$ in Step 5. Finally, according to Remark 2, the selected eigenvalue $\lambda_{k_{t}}^{t}$ is squeezed to $\frac{\lambda_{k_{t}}^{t}\sigma^{2}}{\lambda_{k_{t}}^{t}+\sigma^{2}}$, while the other eigenvalues of ${\bm{\Sigma}}_{t}$ are preserved, to attain the eigenvalues of the next kernel ${\bm{\Sigma}}_{t+1}$. These steps are repeatedly executed until all observation vectors are generated, which completes the ice-filling algorithm.
IV-C Ice-Filling Versus Water-Filling
Figure 3: Comparison between water-filling and ice-filling. The rank of the prior kernel, the number of antennas, and the total pilot length are set as $K=6$, $M=128$, and $Q=16$, respectively. The eigenvalues of the prior kernel, ${\bf\Sigma}_{{\mathbf{h}}}$, are in a descending order, i.e., $\lambda_{1}>\lambda_{2}>\lambda_{3}>\lambda_{4}>\lambda_{5}>\lambda_{6}$.
In this subsection, we put forward a more insightful interpretation to the ice-filling algorithm that establishes its connection with the water-filling principle. Generally speaking, the proposed ice-filling algorithm can be viewed as a quantization of the water-filling principle. Specifically, it is clear that the observation matrices generated by both the water-filling and the ice-filling algorithms fall within the eigenspace of ${\bm{\Sigma}}_{\bf h}$. The major distinction is attributed to their “power allocation strategies”. The water-filling principle allocates $Q$ units of power to the $K$ eigenvectors using (7). The ice-filling algorithm transforms this power-allocation process into an eigenvector-assignment process, where the $K$ eigenvectors are repeatably assigned to the observation vectors in $Q$ timeslots.
To be more precise, recall that $p_{k}$ in (7) denotes the optimal power allocated to the $k$-th eigenvector, satisfying $\sum_{k=1}^{K}p_{k}=Q$. For a clear comparison, we introduce the definition of “pilot reuse frequency” as follows.
Definition 1
The pilot reuse frequency $n_{k}^{t}\in\mathbb{Z}^{+}$, with $k\in\{1,2,\cdots,K\}$ and $t\in\{1,2,\cdots,Q\}$, is defined as the number of times that the $k$-th eigenvector, ${\mathbf{u}}_{k}$, of the prior kernel is selected as the observation vector by the ice-filling algorithm during timeslots $1\sim t$. The pilot reuse frequencies satisfy $\sum_{k=1}^{K}n_{k}^{t}=t$. For ease of expression, we define $n_{k}:=n_{k}^{Q}$ as the pilot reuse frequency during the total $Q$ timeslots.
Comparing the definitions of $p_{k}$ and $n_{k}$, we can intuitively interpret the $n_{k}$-timeslot reuse of the observation vector ${\mathbf{u}}_{k}$ for pilot transmission as allocating $n_{k}$ units of power to ${\mathbf{u}}_{k}$. More importantly, Definition 1 enables us to derive the analytical expressions of the eigenvalues $\{\lambda_{k}^{t}\}_{k=1}^{K}$ in Algorithm 1. Note that the eigenvalue update rule of the ice-filling algorithm can be rewritten as a recursive formula:
$\lambda_{{k_{t}}}^{t+1}=\frac{\lambda_{k_{t}}^{t}\sigma^{2}}{\lambda_{k_{t}}^{t}+\sigma^{2}}\Leftrightarrow\frac{\sigma^{2}}{\lambda_{k_{t}}^{t+1}}=1+\frac{\sigma^{2}}{\lambda_{k_{t}}^{t}},$ (14)
where $\lambda_{k_{t}}^{t}$ is the largest eigenvalue from $\{\lambda_{1}^{t},\lambda_{2}^{t},\cdots,\lambda_{K}^{t}\}$. This recursive formula indicates that whenever the $k$-th eigenvector is selected as an observation vector, the value of $\frac{\sigma^{2}}{\lambda_{k}^{t}}$ increases by 1. By further considering the definition of the pilot reuse frequency $n_{k}^{t}$, that the $k$-th eigenvector is selected by $n_{k}^{t}$ times during timeslots $1\sim t$, we can naturally obtain the relationship between $n_{k}^{t}$ and $\frac{\sigma^{2}}{\lambda_{k}^{t}}$:
$\frac{\sigma^{2}}{\lambda_{k}^{t}}=n_{k}^{t}+\frac{\sigma^{2}}{\lambda_{k}}\Leftrightarrow n_{k}^{t}=\frac{\sigma^{2}}{\lambda_{k}^{t}}-\frac{\sigma^{2}}{\lambda_{k}},\>\>k\in\{1,\cdots,K\}.$ (15)
The comparison between equations (15) and (7) allows us to interpret the ice-filling algorithm as a quantization of the water-filling algorithm, as elaborated below.
Interpretation of ice-filling: As depicted in Fig. 3, the water-filling principle allocates $Q$ units of water (power) to a vessel with $K$ channels, each having a unique base level $\frac{\sigma^{2}}{\lambda_{k}}$, $k\in\{1,2,\cdots,K\}$. By controlling the uniform water level $\beta$, the optimal power $p_{k}$ is determined by the gap between the base level and the water level. In contrast, the ice-filling algorithm transforms the total $Q$ units of water into $Q$ ice blocks, each containing one unit of power and being used for one pilot transmission. Our algorithm starts from an empty vessel and fills one ice block onto the channel having the deepest base surface $\frac{\sigma^{2}}{\lambda_{k_{0}}}$. This operation is equivalent to finding the largest eigenvalue $\lambda_{k_{0}}$ from $\{\lambda_{1},\cdots,\lambda_{K}\}$. Then, the eigenvector ${\mathbf{u}}_{k_{0}}$ is assigned to the first observation vector ${\mathbf{w}}_{1}$, and ice level of this channel increases from $\frac{\sigma^{2}}{\lambda_{k_{0}}}$ to $\frac{\sigma^{2}}{\lambda_{k_{0}}^{1}}=\frac{\sigma^{2}}{\lambda_{k_{0}}}+1$. In the subsequent time slots, the remaining $Q-1$ ice blocks are filled onto the channels locating at the deepest base surfaces or ice levels, indexed by $k_{t}=\arg\min_{k}\{\frac{\sigma^{2}}{\lambda_{k}^{t}}\}_{k=1}^{K}$, one by one. The corresponding eigenvectors ${\mathbf{u}}_{k_{t}}$ are used for pilot transmission, and the ice levels increase according to (15). Consequently, the final pilot reuse frequencies $\{n_{k}\}_{k=1}^{K}$ are determined by the number of ice blocks on top of each channel, and the ice levels are given by $\{\frac{\sigma^{2}}{\lambda_{k}^{Q}}\}_{k=1}^{K}=\{n_{k}+\frac{\sigma^{2}}{\lambda_{k}}\}_{k=1}^{K}$.
The comparison between the ice-filling and water-filling in Fig. 3 indicates that the continuous powers $\{p_{k}\}_{k=1}^{K}$ are quantized into the discrete pilot reuse frequencies $\{n_{k}\}_{k=1}^{K}$. This quantization results in the non-uniform ice levels $\{\frac{\sigma^{2}}{\lambda_{k}^{Q}}\}_{k=1}^{K}$, as opposed to the uniform water level $\beta$. To see the quantization nature of ice-filling more clearly, we prove Theorem 2 to upper bound the quantization error between $p_{k}$ and $n_{k}$.
Theorem 2
Considering the $k$-th pilot reuse frequency, $n_{k}$, and the $k$-th optimal power, $p_{k}$, we have
$|n_{k}-p_{k}|<1.$ (16)
Proof:
See Appendix B. ∎
Remark 3
Theorem 2 rigorously proves that the deviation of $n_{k}$ from $p_{k}$ is less than 1, rendering the pilot reuse frequency a good approximation of the optimal power allocation. Particularly, when the total pilot length $Q$ is large, the relative error between $n_{k}$ and $p_{k}$ tends to zero because
$\frac{|n_{k}-p_{k}|}{p_{k}}<\frac{1}{p_{k}}\overset{Q\rightarrow+\infty}{=}0.$ (17)
Thereafter, the proposed ice-filling algorithm features a discrete approximation to the ideal water-filling algorithm with the relative quantization error decaying rapidly.
IV-D Majorization Minimization for Phase-Only Controllable Combiner
Turn now to the phase-only controllable analog combiners where the amplitude of each element of $\bf W$ is fixed as a constant. The observation vector at the $(t+1)$-th timeslot can be obtained by
${{\bf{w}}_{t+1}}=\mathop{\arg\!\max}\limits_{\left|{\bf{w}}\right|=\frac{1}{{\sqrt{M}}}{{\bf{1}}_{M}}}{{\bf{w}}^{H}}{{\bf{\Sigma}}_{t}}{\bf{w}}.$ (18)
Due to the unit modulus constraint, the problem (18) is a non-convex programming. To cope with this problem, we propose a MM-based observation matrix design, as presented in Algorithm 2. Its key idea is to reformulate the non-convex problem as a series of more tractable approximate subproblems, which can be solved in an alternating manner [50]. Specifically, the proposed MM-based method is illustrated as follows.
Algorithm 2 MM-Based Observation Matrix Design
0:  Number of pilots $Q$, kernel ${\bm{\Sigma}}_{\bf h}$.
0:  Designed observation matrix ${\bf W}$.
1:  Initialization: ${\bf W}=\varnothing$, ${{\bf{\Sigma}}_{0}}={\bm{\Sigma}}_{\bf h}$.
2:  for $t=0,\cdots,Q-1$ do
3:     Initialize ${\bf w}_{t+1}$ randomly
4:     Update: ${\bf{X}}=\mathsf{Tr}\left({{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}-{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}$
5:     while no convergence of ${{\bf{w}}^{H}_{t+1}}{{\bf{\Sigma}}_{t}}{\bf{w}}_{t+1}$ do
6:        Update: ${\bf v}={\bf w}_{t+1}$
7:        Update the $t+1$-th observation vector: ${\bf{w}}_{t+1}=\frac{1}{{\sqrt{M}}}{\exp\left(j\angle\left({{\bf{X}}-{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}+{{\bf{\Sigma}}_{t}}}\right){\bf{v}}\right)}$
8:     end while
9:     Update kernel: ${\bf{\Sigma}}_{t+1}={\bf{\Sigma}}_{t}-\frac{{\bf{\Sigma}}_{t}{\bf{w}}_{t+1}{\bf{w}}_{t+1}^{H}{\bf{\Sigma}}_{t}}{{\bf{w}}_{t+1}^{H}{\bf{\Sigma}}_{t}{\bf{w}}_{t+1}+\sigma^{2}}$
10:  end for
11:  Construct observation matrix: ${\bf W}=[{\mathbf{w}}_{1},{\mathbf{w}}_{2},\cdots,{\mathbf{w}}_{Q}]$
12:  return  Designed observation matrix ${\bf W}$
To begin with, given that adding a constant to the objective function does not affect the optimal solution of (18), problem (18) is equivalent to
${{\bf{w}}_{t+1}}\overset{(a)}{=}$ $\mathop{\arg\!\max}\limits_{\left|{\bf{w}}\right|=\frac{1}{{\sqrt{M}}}{{\bf{1}}_{M}}}{{\bf{w}}^{H}}{{\bf{\Sigma}}_{t}}{\bf{w}}-{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{w}}^{H}}{\bf{w}}$
$=$ $\mathop{\arg\!\min}\limits_{\left|{\bf{w}}\right|=\frac{1}{{\sqrt{M}}}{{\bf{1}}_{M}}}{{\bf{w}}^{H}}\left({{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}-{{\bf{\Sigma}}_{t}}}\right){\bf{w}},$ (19)
where (a) holds because ${\mathbf{w}}^{H}{\mathbf{w}}=1$. Then, we consider to solve problem (IV-D) by iteratively minimizing the upper-bound of its objective function. To this end, we introduce an auxiliary variable ${\bf v}\in{\mathbb{C}}^{M\times 1}$ to obtain an upper-bound function ${\bar{f}}({\bf w},{\bf v})$, given by [51]:
${{\bf{w}}^{H}}$ $\left({{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}-{{\bf{\Sigma}}_{t}}}\right){\bf{w}}\leq{\bar{f}}({\bf w},{\bf v})=$
${{\bf{w}}^{H}}{\bf{Xw}}-2\Re\left\{{{{\bf{w}}^{H}}\left({{\bf{X}}-{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}+{{\bf{\Sigma}}_{t}}}\right){\bf{v}}}\right\}+$
${{\bf{v}}^{H}}\left({{\bf{X}}-{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}+{{\bf{\Sigma}}_{t}}}\right){\bf{v}},$ (20)
where ${\bf{X}}=\mathsf{Tr}\left({{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}-{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}$ and the equality holds when ${\bf v}={\bf w}$. In this way, ${\bf w}$ and ${\bf v}$ in the upper-bound function ${\bar{f}}({\bf w},{\bf v})$ can be alternatively optimized to approach a sub-optimal solution to problem (IV-D) [50]. In particular, for a given ${\bf w}$, ${\bf v}$ can be updated by ${\bf v}={\bf w}$. For a given ${\bf v}$, by utilizing ${{\bf{w}}^{H}}{\bf{X}}{\bf{w}}=\mathsf{Tr}\left({{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}-{{\bf{\Sigma}}_{t}}}\right)$ and removing the unrelated part in ${\bar{f}}({\bf w},{\bf v})$ in (IV-D), the subproblem of updating $\bf w$ can be rewritten as
${{\bf{w}}}_{t+1}=$
$\mathop{\arg\!\max}\limits_{\left|{\bf{w}}\right|=\frac{1}{{\sqrt{M}}}{{\bf{1}}_{M}}}\Re\left\{{{{\bf{w}}^{H}}\left({{\bf{X}}-{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}+{{\bf{\Sigma}}_{t}}}\right){\bf{v}}}\right\}.$ (21)
It can be easily proved that the optimal solution to (IV-D) is given by ${{\bf{w}}}_{t+1}=\frac{1}{{\sqrt{M}}}{\exp\left(j\angle\left({{\bf{X}}-{\lambda_{\max}}\left({{{\bf{\Sigma}}_{t}}}\right){{\bf{I}}_{M}}+{{\bf{\Sigma}}_{t}}}\right){\bf{v}}\right)}$. This solution, associated with the updating approach of the posterior kernel in (9), completes the algorithm. It is worth noting that, since the updates of $\bf v$ and ${\bf w}_{t+1}$ both monotonously increase the objective function, the convergence of Algorithm 2 is naturally guaranteed.
IV-E Kernel Selection in Non-Ideal Case
In some non-ideal cases where the perfect prior kernel, ${\bf\Sigma}_{\bf h}$, is difficult to acquire, the initialization states of Algorithms 1 and 2 may pose a challenge. To deal with this issue, we suggest two methods to obtain imperfect prior kernels.
IV-E1 Statistical kernel
Firstly, the prior kernel can be selected as the statistical covariance of historical estimated channels, i.e., ${\bf{\hat{\Sigma}}}_{{\bf h}}=\mathsf{E}({{\bf{\hat{h}}}{{{\bf{\hat{h}}}}^{H}}})$. Due to the error of channel estimation, the historical channels ${\bf{\hat{h}}}$ deviate from the true channels ${\bf{h}}$. We use a Gaussian noise to model this deviation, and get ${\bf{\hat{h}}}={\bf{h}}+{{\bf{z}}_{\bf{h}}}$, wherein ${{\bf{z}}_{\bf{h}}}\sim{\cal CN}\left({{{\bf{0}}_{M}},\sigma_{\bf{h}}^{2}{{\bf{I}}_{M}}}\right)$ denotes the historical channel estimation error. For simplicity, we name $\sigma_{\bf{h}}^{2}$ as the kernel estimation error. Thereby, the statistical kernel is modeled as
${\bf{\hat{\Sigma}}}_{{\bf h}}={\bf{\Sigma}}_{{\bf h}}+\sigma^{2}_{{\mathbf{h}}}{\mathbf{I}}_{M}.$ (22)
IV-E2 Artificial kernels
Another choice is to employ artificially designed kernels to mimic the perfect kernel. To this end, the artificial kernels should assign high similarity to nearby antennas while diminishing the correlation rapidly with inter-antenna distance. For example, here we consider a general uniform planer array (UPA) equipped with $M$ antennas with an antenna spacing of $d$. The numbers of its horizontal antennas and vertical antennas are $M_{x}$ and $M_{y}$, respectively. Striking a balance between complexity and practicality, we recommend two artificial kernels as follows.
  • Exponential kernel: The exponential kernel ${\bm{\Sigma}}_{\rm exp}$, is the most popular selection in Bayesian estimation. Let ${{\bf{m}}_{x}}=\left[{-\frac{{{M_{x}}-1}}{2},-\frac{{{M_{x}}-3}}{2},\cdots,\frac{{{M_{x}}-1}}{2}}\right]^{T}$ and ${{\bf{m}}_{y}}=\left[{-\frac{{{M_{y}}-1}}{2},-\frac{{{M_{y}}-3}}{2},\cdots,\frac{{{M_{y}}-1}}{2}}\right]^{T}$. Then, the exponential kernels for the two dimensions of UPA, ${{\bf{\Sigma}}_{{\rm exp},x}}\in{\mathbb{C}}^{M_{x}\times M_{x}}$ and ${{\bf{\Sigma}}_{{\rm exp},y}}\in{\mathbb{C}}^{M_{y}\times M_{y}}$, can be respectively expressed as
    ${{\bf{\Sigma}}_{{\rm exp},x}}\!=\!\exp\left({\!-{\eta_{1}^{2}}{\frac{4\pi^{2}d^{2}}{\lambda^{2}}}{{\left|{{{\bf{1}}^{T}_{{M_{x}}}}\otimes{{\bf{m}}_{x}}\!-\!{\bf{m}}_{x}^{T}\otimes{{\bf{1}}_{{M_{x}}}}}\right|}^{\odot 2}}}\right),$ (23)
    ${{\bf{\Sigma}}_{{\rm exp},y}}\!=\!\exp\left({\!-{\eta_{1}^{2}}{\frac{4\pi^{2}d^{2}}{\lambda^{2}}}{{\left|{{{\bf{1}}^{T}_{{M_{y}}}}\otimes{{\bf{m}}_{y}}\!-\!{\bf{m}}_{y}^{T}\otimes{{\bf{1}}_{{M_{y}}}}}\right|}^{\odot 2}}}\right),$ (24)
    where $\eta_{1}$ is an adjustable hyper-parameter and ${\bf X}^{\odot 2}$ denotes the element-wise product of two matrices $\bf X$. Then, the overall kernel can be written as
    ${\bm{\Sigma}}_{\rm exp}={{\bf{\Sigma}}_{{\rm exp},x}}\otimes{{\bf{\Sigma}}_{{\rm exp},y}}.$ (25)
    The exponential kernel, ${\bm{\Sigma}}_{\rm exp}$, is a heuristic kernel. It exhibits the highest correlation at the diagonal elements, with the channel correlation diminishing exponentially as the antenna spacing increases. Compared to other kernels, the exponential kernel shows reduced sensitivity to outliers, making it suitable for reconstructing channels that lack obvious regularity.
  • Bessel kernel: The Bessel kernel, denoted as ${\bm{\Sigma}}_{\rm bes}$, is well-suited for capturing and modeling complex-valued data exhibiting oscillatory patterns. It characterizes the covariance of array steering vector when the user distributes uniformly in the spatial domain. For the considered UPA, the Bessel kernel can be obtained as ${\bm{\Sigma}}_{\rm bes}=\mathsf{E}({\mathbf{a}}(\theta,\phi){\mathbf{a}}^{H}(\theta,\phi))$, where ${\mathbf{a}}(\theta,\phi)$ is the array steering vector of UPA, and $\theta\sim\mathcal{U}(-\pi/2,\pi/2)$ and $\phi\sim\mathcal{U}(-\pi/2,\pi/2)$ denote the angle-of-arrivals. By calculating expectation, the Bessel kernel is expressed as
    ${{\bf{\Sigma}}_{{\rm{bes}},x}}$ $={J_{0}}\left({\eta_{2}\frac{2\pi d}{\lambda}\left|{{\bf{1}}_{{M_{x}}}^{T}\otimes{{\bf{m}}_{x}}-{\bf{m}}_{x}^{T}\otimes{{\bf{1}}_{{M_{x}}}}}\right|}\right),$ (26)
    ${{\bf{\Sigma}}_{{\rm{bes}},y}}$ $={J_{0}}\left({\eta_{2}\frac{2\pi d}{\lambda}\left|{{\bf{1}}_{{M_{y}}}^{T}\otimes{{\bf{m}}_{y}}-{\bf{m}}_{y}^{T}\otimes{{\bf{1}}_{{M_{y}}}}}\right|}\right),$ (27)
    where $J_{0}$ is the zero-order Bessel function of the first kind and $\eta_{2}$ is a hyper-parameter. Thus, the overall kernel can be written as
    ${{\bf{\Sigma}}_{{\rm{bes}}}}={{\bf{\Sigma}}_{{\rm{bes}},x}}\otimes{{\bf{\Sigma}}_{{\rm{bes}},y}}.$ (28)
V Performance Analysis
To evaluate the performance of the proposed MIM-based observation matrix design, this section provides comprehensive MSE analyses under different algorithmic conditions.
V-A Estimation Accuracy Under Perfect Kernel
When the perfect prior kernel ${{\bf{\Sigma}}}_{{\mathbf{h}}}$ is available, the MSEs achieved by the proposed algorithms can be determined by the trace of their posterior kernels, i.e.,
$\mathsf{E}(\|{\bm{\mu}_{{\mathbf{h}}|{\mathbf{y}}}}-{\mathbf{h}}\|^{2}_{2})=\mathsf{Tr}\left({{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}\right).$ (29)
The subsequent two lemmas give out the close-form MSEs and their asymptotic expressions of the water-filling-based and ice-filling-based channel estimators, respectively.
Lemma 1
If the perfect kernel ${{\bf{\Sigma}}}_{{\mathbf{h}}}$ is adopted, the MSE $\delta_{\rm wf}$ of the water-filling algorithm is given by
$\delta_{\rm wf}=\sum_{k=1}^{K}\frac{\lambda_{k}\sigma^{2}}{p_{k}\lambda_{k}+\sigma^{2}}\overset{Q\rightarrow+\infty}{=}\mathcal{O}(K^{2}Q^{-1}).$ (30)
Proof:
See Appendix C. ∎
Lemma 2
If the perfect kernel ${{\bf{\Sigma}}}_{{\mathbf{h}}}$ is adopted, the MSE $\delta_{\rm if}$ of the ice-filling algorithm is given by
$\delta_{\rm if}=\sum_{k=1}^{K}\frac{\lambda_{k}\sigma^{2}}{n_{k}\lambda_{k}+\sigma^{2}}\overset{Q\rightarrow+\infty}{=}\mathcal{O}(K^{2}Q^{-1}).$ (31)
Proof:
See Appendix D. ∎
We can gain two insights from Lemma 1 and Lemma 2. Firstly, the MSEs achieved by the water-filling and ice-filling methods have identical asymptotic expressions, both of which decay at the rate of $Q^{-1}$. The only difference is owing to the quantization of the continuous powers $\{p_{k}\}$ in (30) to the pilot reuse frequencies $\{n_{k}\}$ in (31). Next, the MSEs decay quadratically as the rank of the channel kernel, $K$, decreases. This finding demonstrates the superiority of DASs in channel estimation. A denser antenna deployment increases the correlation between inter-antenna channels, decreases the rank of the channel kernel, and thereby improves the channel estimation accuracy.
Taking into account the quantization nature revealed in Theorem 2, the asymptotic achievability bound of the MSE of the ice-filling algorithm can be evaluated by the following theorem.
Theorem 3
As $Q\rightarrow+\infty$, the asymptotic MSE difference, $|\delta_{\rm wf}-\delta_{\rm if}|$, is given by
$|\delta_{\rm wf}-\delta_{\rm if}|=\mathcal{O}(K^{3}Q^{-2}).$ (32)
Proof:
See Appendix E. ∎
Remark 4
Theorem 3 guarantees that the MSE gap between the ice-filling and the ideal water-filling algorithms decays at a rate of $K^{3}Q^{-2}$, which demonstrates the near-optimality of the proposed ice-filling channel estimator when $Q\rightarrow+\infty$ or $K\rightarrow 0$.
To validate the superiority of the proposed design, it is of great interest to evaluate the MSE achieved by the Bayesian regression (4) using randomly generated observation matrices. For amplitude-and-phase controllable cases, we assume that all elements of ${\mathbf{W}}$ are independently generated from a Gaussian distribution $\mathcal{CN}(0,1/M)$. For phase-only controllable cases, the angles of all elements of $\bf W$ are randomly selected from $[-\pi,+\pi]$. Then, we derive the asymptotic MSE of a random observation matrix in Lemma 3.
Lemma 3
Assume that the perfect kernel ${{\bf{\Sigma}}}_{{\mathbf{h}}}$ is adopted and $Q$ is sufficiently large. When $\bf W$ is randomly generated, the MSE $\delta_{\rm rnd}$ achieved in both phase-and-amplitude and phase-only controllable cases can be approximated by
$\delta_{\rm rnd}\approx\sum\limits_{k=1}^{K}{{{{\lambda_{k}}{\sigma^{2}}}\over{{Q\over M}{\lambda_{k}}+{\sigma^{2}}}}}\overset{Q\rightarrow+\infty}{=}\mathcal{O}(KMQ^{-1}),$ (33)
where the approximate equality can be infinitely close to an equality as $Q\to\infty$.
Proof:
See Appendix F. ∎
It is evident from Lemmas 1-3 that meticulously designed observation matrices can scale down the asymptotic MSE from $\mathcal{O}(KMQ^{-1})$ to $\mathcal{O}(K^{2}Q^{-1})$. This improvement is attributed to the fact that the observation matrices designed by water-filling and ice-filling approaches are tightly aligned with the kernel’s eigenspace of dimension $K$. In contrast, the random observation matrix wastes a large amount of power in the kernel’s null space of dimension $M-K$, which contains no useful information about the channel. This finding demonstrates the significant role played by the observation matrix to enhance channel estimation accuracy, particularly in DASs where $M\gg K$.
V-B Estimation Accuracy Under Imperfect Kernel
We now consider the situation where the perfect prior kernel ${\bm{\Sigma}}_{\bf h}$ is not available. In this context, an imperfect prior kernel, such as the statistical kernel in (22) and the artificial kernels (25) and (28), has to be adopted in the proposed algorithms. We denote the adopted prior kernel as $\bf\Sigma$. Then, one can use the following lemma to evaluate the MSE performance analytically.
Lemma 4
Given a prior kernel $\bf\Sigma$ as the input of the proposed channel estimator, the MSE $\hat{\delta}$ of channel estimation is given by
${\hat{\delta}}=\mathsf{Tr}\left(\left({{{\bf{\Pi}}^{H}}{{\bf{W}}^{H}}-{{\bf{I}}_{M}}}\right){{\bf{\Sigma}}_{\bf{h}}}\!\left({{\bf{W\Pi}}-{{\bf{I}}_{M}}}\right)\right)+{\sigma^{2}}\mathsf{Tr}\!\left({{\bf{\Pi}}^{H}}{\bf{\Pi}}\right),$ (34)
where ${\bf{\Pi}}:={\left({{{\bf{W}}^{H}}{\bf{\Sigma W}}+{\sigma^{2}}{\bf I}_{Q}}\right)^{-1}}{{\bf{W}}^{H}}{\bf{\Sigma}}$.
Proof:
See Appendix G. ∎
Lemma 4 indicates that, the estimation error caused by the non-ideal input cannot be neglected. For example, if the input kernel $\bm{\Sigma}$ behaves far from the real kernel ${\bm{\Sigma}}_{\bf h}$, the value of the MSE may become significantly large due to the existence of $\bm{\Pi}$ in (34). To obtain more insightful results, we use the statistical kernel, ${\bf\hat{\Sigma}}_{{\mathbf{h}}}$ in (22), as an example to analyze the achievable MSEs. To ease the derivation, we first prove a useful expression of the MSE in Lemma 5.
Lemma 5
The MSE $\hat{\delta}$ in (34) under the imperfect kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}$ can be written as a function of ${{\bf{W}}{{\bf{W}}^{H}}}$, given by
$\hat{\delta}=$ $\mathsf{Tr}\left({\left({{\bf{\hat{\Sigma}}}_{\bf{h}}^{H}{{\bf{\Omega}}^{H}}-{{\bf{I}}_{M}}}\right){{\bf{\Sigma}}_{\bf{h}}}\left({{\bf{\Omega}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}-{{\bf{I}}_{M}}}\right)}\right)$
$+{\sigma^{2}}\mathsf{Tr}\left({{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\bf{\Xi}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}}\right),$ (35)
wherein ${\bf{\Omega}}$ and ${\bf{\Xi}}$ are subfunctions of ${{\bf{W}}{{\bf{W}}^{H}}}$, written as
${\bf{\Omega}}=$ ${{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{2}}}}-{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{4}}}}{\left({{\bf{\hat{\Sigma}}}_{\bf{h}}^{-1}+{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{2}}}}}\right)^{-1}}{\bf{W}}{{\bf{W}}^{H}},$ (36)
${\bf{\Xi}}=$ ${{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{4}}}}-2{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{6}}}}{\left({{\bf{\hat{\Sigma}}}_{\bf{h}}^{-1}+{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{2}}}}}\right)^{-1}}{\bf{W}}{{\bf{W}}^{H}}+$
${{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{8}}}}\left({\left({{\bf{\hat{\Sigma}}}_{\bf{h}}^{-1}+{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{2}}}}}\right)^{-1}}{\bf{W}}{{\bf{W}}^{H}}\right)^{2}.$ (37)
Proof:
See Appendix H. ∎
Recall that the essential difference among the three estimators, i.e., water-filling algorithm, ice-filling algorithm, and the estimator with randomly generated $\bf W$, is the form of observation matrix $\bf W$. Thus, by utilizing Lemma 5, the MSE of different algorithms under the imperfect kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}$ can be obtained by replacing ${{\bf{W}}{{\bf{W}}^{H}}}$ in $\hat{\delta}$ with their corresponding analytical expressions. Aided by some matrix operations, we can thereby prove the following lemma.
Lemma 6
Under the imperfect kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}$, the achievable MSE for the water-filling algorithm can be written as
${{\hat{\delta}}_{{\rm{wf}}}}$ $={\sigma^{2}}\sum\limits_{k=1}^{K}{\frac{{{\lambda_{k}}{\sigma^{2}}\!+\!{{\hat{p}}_{k}}{{\left({{\lambda_{k}}\!+\!\sigma_{\bf{h}}^{2}}\right)}^{2}}}}{{{{\left({{\hat{p}_{k}}\left({\lambda_{k}\!+\!\sigma_{\bf{h}}^{2}}\right)+{\sigma^{2}}}\right)}^{2}}}}}\!+\!\sigma^{2}\!\!\!\!\sum\limits_{m=K+1}^{M}{\frac{{{{\hat{p}}_{m}}\sigma_{\bf{h}}^{4}}}{{{{\left({{{\hat{p}}_{m}}\sigma_{\bf{h}}^{2}\!+\!{\sigma^{2}}}\right)}^{2}}}}}$
$\mathop{=}\limits^{Q\to+\infty}{\cal O}\left({{\sigma^{2}}{M^{2}}{Q^{-1}}}\right),$ (38)
wherein $\{{\lambda}_{1},\cdots,{\lambda}_{M}\}$ represent the $M$ non-negative eigenvalues of ${\bm{\Sigma}}_{\bf h}$ in a descending order [^1] That is to say, if ${\bm{\Sigma}}_{\bf h}$ is rank-$K$, we have ${\lambda}_{m}=0$ for all $m>K$.. The power ${\hat{p}}_{m}$ allocated to each eigenvector is calculated by the water-filling principle ${\hat{p}}_{m}=\left({\hat{\beta}}-\frac{\sigma^{2}}{\lambda_{m}+\sigma^{2}_{\bf h}}\right)^{+}$ for $m\in\{1,\cdots,M\}$. The water-level ${\hat{\beta}}$ can be determined via binary search such that $\sum_{m=1}^{M}{\hat{p}}_{m}=Q$.
Under the imperfect kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}$, the achievable MSE for the ice-filling algorithm can be written as
${{\hat{\delta}}_{{\rm{if}}}}$ $={\sigma^{2}}\sum\limits_{k=1}^{K}{\frac{{{\lambda_{k}}{\sigma^{2}}\!+\!{{\hat{n}}_{k}}{{\left({{\lambda_{k}}\!+\!\sigma_{\bf{h}}^{2}}\right)}^{2}}}}{{{{\left({{\hat{n}_{k}}\left({\lambda_{k}\!+\!\sigma_{\bf{h}}^{2}}\right)\!+\!{\sigma^{2}}}\right)}^{2}}}}}\!+\!\sigma^{2}\!\!\!\!\sum\limits_{m=K+1}^{M}{\frac{{{{\hat{n}}_{m}}\sigma_{\bf{h}}^{4}}}{{{{\left({{{\hat{n}}_{m}}\sigma_{\bf{h}}^{2}\!+\!{\sigma^{2}}}\right)}^{2}}}}}$
$\mathop{=}\limits^{Q\to+\infty}{\cal O}\left({{\sigma^{2}}{M^{2}}{Q^{-1}}}\right),$ (39)
wherein the pilot reuse frequencies $\{{\hat{n}}_{1},\cdots,{\hat{n}}_{M}\}$ can be obtained by Algorithm 1 with ${\bf{\hat{\Sigma}}}_{{\bf h}}$ being the input kernel. In particular, $\sum_{m=1}^{M}{\hat{n}}_{m}=Q$ holds.
For the estimator enabled by a randomly generated $\bf W$ with sufficiently large $Q$, the achievable MSE under the imperfect kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}$ can be written as
${{\hat{\delta}}_{{\rm{rnd}}}}$ $\approx{\sigma^{2}}\sum\limits_{k=1}^{K}{\frac{{{\lambda_{k}}{\sigma^{2}}\!+\!\frac{Q}{M}{{\left({{\lambda_{k}}\!+\!\sigma_{\bf{h}}^{2}}\right)}^{2}}}}{{{{\left({\frac{Q}{M}\left({\lambda_{k}\!+\!\sigma_{\bf{h}}^{2}}\right)\!+\!{\sigma^{2}}}\right)}^{2}}}}}\!+\!\sigma^{2}\!\!\!\!\sum\limits_{m=K+1}^{M}{\frac{{\frac{Q}{M}\sigma_{\bf{h}}^{4}}}{{{{\left({\frac{Q}{M}\sigma_{\bf{h}}^{2}\!+\!{\sigma^{2}}}\right)}^{2}}}}}$
$\mathop{=}\limits^{Q\to+\infty}{\cal O}\left({{\sigma^{2}}{M^{2}}{Q^{-1}}}\right),$ (40)
wherein the approximate equality can be infinitely close to an equality as $Q\to\infty$.
Proof:
See Appendix I. ∎
From Lemma 6, it is easy to prove that the MSEs under the imperfect kernel $\hat{\bf\Sigma}_{\bf h}$ are increased by the kernel estimation error $\sigma_{\bf h}^{2}$. The reason is that, the estimation error makes the input kernel full-rank, i.e., ${\hat{\bf{\Sigma}}_{\bf{h}}}={{\bf{\Sigma}}_{\bf{h}}}+\sigma_{\bf{h}}^{2}{{\bf{I}}_{M}}$. Then, the power (or pilot reuse frequency) for estimation may be allocated to all eigenvectors of ${\bf\Sigma}_{\bf h}$. However, allocating power to those eigenvectors associated with the eigenvalue $\sigma_{\bf{h}}^{2}$ contributes nearly no improvement to channel recovery accuracy. Besides, an imperfect kernel input also changes the weights of the posterior mean ${\bm{\mu}}_{{\bf h}|{\bf y}}$ in (4), which resembles a Bayesian estimator that underestimates the impact of noise.
Remark 5
According to (6), (6), and (6), an efficient way to alleviate the effect of $\sigma^{2}_{\bf h}$ on MSEs is to introduce the prior knowledge that the real kernel $\bf{\Sigma}_{\bf h}$ is rank-$K$. Then, the estimator only allocates power to the eigenvalues associated with the $K$-largest eigenvalues of ${\hat{\bf{\Sigma}}_{\bf{h}}}$, so that the second term in (6), (6), and (6) can be eliminated. For example, when $p_{m}$ for $m>K$ are forced to be zero, ${{\hat{\delta}}_{{\rm{wf}}}}$ in (6) becomes
${\hat{\delta}}_{{\rm{wf}}}$ $={\sigma^{2}}\sum\limits_{k=1}^{K}{{{{\lambda_{k}}{\sigma^{2}}+{p_{k}}{{\left({{\lambda_{k}}+\sigma_{\bf{h}}^{2}}\right)}^{2}}}\over{{{\left({{p_{k}}\left({{\lambda_{k}}+\sigma_{\bf{h}}^{2}}\right)+{\sigma^{2}}}\right)}^{2}}}}}$
$\mathop{=}\limits^{Q\to+\infty}{\cal O}\left({{\sigma^{2}}{K^{2}}{Q^{-1}}}\right),$ (41)
wherein $p_{k}$ for $k\in\{1,\cdots,K\}$ is given in (7). For $\sigma_{\bf{h}}^{2}>0$, it is evident that ${\hat{\delta}}_{{\rm{wf}}}$ in (5) is lower than that in (6).
Based on the results in Lemma 6, the asymptotic MSEs when the kernel estimation error $\sigma_{{\mathbf{h}}}^{2}$ is infinitely large are expressed as follows.
Corollary 1
When $\sigma_{\bf{h}}^{2}\to+\infty$, the asymptotic MSEs ${\hat{\delta}}$ for the water-filling algorithm, ice-filling algorithm, and the estimator based on random $\bf W$ can be respectively written as
${\hat{\delta}}_{\rm wf}$ $\mathop{=}\limits^{\sigma_{\bf{h}}^{2}\to+\infty}{\cal O}\left({{\sigma^{2}}\sum\nolimits_{k=1}^{{L_{{\rm{wf}}}}}{\hat{p}_{k}^{-1}}}\right)={\cal O}\left({{\sigma^{2}}{M^{2}}{Q^{-1}}}\right),$ (42)
${\hat{\delta}}_{\rm if}$ $\mathop{=}\limits^{\sigma_{\bf{h}}^{2}\to+\infty}{\cal O}\left({{\sigma^{2}}\sum\nolimits_{k=1}^{{L_{{\rm{if}}}}}{\hat{n}_{k}^{-1}}}\right)={\cal O}\left({{\sigma^{2}}{M^{2}}{Q^{-1}}}\right),$ (43)
${\hat{\delta}}_{\rm rnd}$ $\mathop{=}\limits^{\sigma_{\bf{h}}^{2}\to+\infty}{\cal O}\left({{\sigma^{2}}{M^{2}}{Q^{-1}}}\right),$ (44)
where $L_{\rm wf}$ and $L_{\rm if}$ are the numbers of the positive values in $\{{\hat{p}}_{m}\}_{m=1}^{M}$ and $\{{\hat{n}}_{m}\}_{m=1}^{M}$, respectively.
Proof:
The three asymptotic MSEs can be easily obtained by letting $\sigma_{\bf{h}}^{2}\to+\infty$ for ${\hat{\delta}}_{\rm wf}$ in (6), ${\hat{\delta}}_{\rm if}$ in (6), and ${\hat{\delta}}_{\rm rnd}$ in (6), respectively. Since their derivations are similar, here we focus on the asymptotic ${\hat{\delta}}_{\rm wf}$ as an example. By letting $\sigma_{\bf{h}}^{2}\to+\infty$ and removing the small-order components, we have ${\hat{\delta}}_{\rm wf}\mathop{=}\limits^{\sigma_{\bf{h}}^{2}\to+\infty}{\cal O}\left({{\sigma^{2}}\sum\nolimits_{k=1}^{{L_{{\rm{wf}}}}}{\hat{p}_{k}^{-1}}}\right)$. Since ${\hat{p}}_{m}=\left({\hat{\beta}}-\frac{\sigma^{2}}{\lambda_{m}+\sigma^{2}_{\bf h}}\right)^{+}$ for $m\in\{1,\cdots,M\}$, we have ${\hat{p}}_{m}={\hat{\beta}}=\frac{Q}{M}$ and $L_{{\rm{wf}}}=M$, which completes the proof. ∎
V-C Computational Complexity Analysis
The computational complexity of the proposed channel estimators is low in practical applications. Specifically, according to Fig. 2, the computational complexity can be divided into two components, i.e., observation matrix design in Stage 1 and Bayesian regression in Stage 2.
For the observation matrix design, the complexity of Algorithm 1 is mainly caused by the eigenvalue decomposition and the principal eigenvalue selection, which is ${\cal O}(M^{3}+QK)$. The complexity of Algorithm 2 mainly depends on the search for the largest eigenvalue and the iterations required for the convergence of MM, which is ${\cal O}(M^{3}+I_{o}M^{2}Q)$ wherein $I_{o}$ is the number of iterations.
For the Bayesian regression, according to (4), we find that the recovered channel $\hat{\bf h}$ is exactly the weighted sum of received pilots $\bf y$. In particular, the weight is calculated by ${\bf{\Sigma}_{\bf{h}}}{{\bf{W}}}{\left({\bf W}^{H}{\bf{\Sigma}}_{\bf{h}}{\bf W}+\sigma^{2}{\bf I}_{Q}\right)^{-1}}$, thus the overall complexity of Stage 2 is ${\cal O}(M^{3})$. It is notable that the weight only relies on the designed ${\bf W}$ and $\bm{\Sigma}_{\bf h}$. It means that the weight for recovering $\bf h$ can also be calculated offline and then employed for online channel estimation.
VI Simulation Results
TABLE I: Simulation Parameters of Channel Model in 3GPP TR 38.901
Channel parameters Values in [52]
Carrier frequency $f_{c}$ 3.5 GHz
Number of clusters 23
Number of rays per cluster 20
Path gains ${\cal CN}(0,1)$
Incident angles ${\cal U}(-90^{\circ},+90^{\circ})$
Maximum angle spread ${\cal U}(-5^{\circ},+5^{\circ})$
Maximum delay spread ${\cal U}(-30~{}{\rm ns},+30~{}{\rm ns})$
In this section, we present simulation results to verify the effectiveness of the proposed channel estimators for DASs.
VI-A Simulation Setup and Benchmarks
To account for a practical scenario, the 3GPP TR 38.901 channel model is used for simulations, whose key parameters are given in Table I. Otherwise specifically specified, we consider a UPA. The number of antennas is set to $M=128$, and the numbers of horizontal antennas and vertical antennas are set to $M_{x}=16$ and $M_{y}=8$, respectively. The antenna spacing is set to $\frac{\lambda}{8}$. The signal-to-noise ratio (SNR) is defined as ${\rm SNR}=\frac{{\mathsf{E}}\left(\|{\bf h}\|^{2}\right)}{\sigma^{2}}$. The performance is evaluated by the normalized mean square error (NMSE), which is defined as ${\rm NMSE}={\mathsf{E}}\left(\frac{\|{\bf h}-\hat{\bf h}\|^{2}}{\|{\bf h}\|^{2}}\right)$. 3000 Monte-Carlo experiments are carried out to plot each figure.
To demonstrate the effectiveness of the proposed ice-filling and MM algorithms, six observation matrices are considered for comparison:
  • Random matrix ${\mathbf{W}}^{\rm Rand}\in\mathbb{C}^{M\times Q}$: The coefficients of ${\mathbf{W}}^{\rm Rand}$ are independently generated from the standard complex Gaussian distribution. Then, each column is normalized to satisfy the unit power constraint.
  • Top-$Q$ matrix ${\mathbf{W}}^{{\rm Top}Q}\in\mathbb{C}^{M\times Q}$: The first $Q$ major eigenvectors of ${\bf\Sigma}_{{\mathbf{h}}}$ are used to set the observation matrix ${\mathbf{W}}^{{\rm Top}Q}$.
  • Ice-filling matrix ${\mathbf{W}}^{\rm IF}\in\mathbb{C}^{M\times Q}$: The proposed ice-filling algorithm is used to generate ${\mathbf{W}}^{\rm IF}$.
  • MM matrix ${\mathbf{W}}^{\rm MM}\in\mathbb{C}^{M\times Q}$: The proposed MM algorithm is used to generate ${\mathbf{W}}^{\rm MM}$.
  • Water-filling matrix ${\mathbf{W}}^{\rm WF}\in\mathbb{C}^{M\times Q}$: The water-filling scheme is employed to generate ${\mathbf{W}}^{\rm WF}$. To implement this ideal observation matrix, the noise vector, ${\mathbf{z}}$, in (2) is assumed to be independent of the observation matrix ${\mathbf{W}}^{\rm WF}$, and its distribution is fixed as $\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$.
  • DFT matrix ${\mathbf{W}}^{\rm DFT}\in\mathbb{C}^{M\times M}$: The DFT matrix requires the pilot length, $Q$, to be equal to the number of antennas, $M$. This matrix is used for performing LS algorithm.
Note that the pilot length for the DFT matrix ${\mathbf{W}}^{\rm DFT}$ is $Q=M=128$, while that for the other matrices is $Q=64$ by default. The above observation matrices are used in four channel estimators for comparison: The LS, vector approximate message passing (VAMP) [28], OMP, and MMSE algorithms. For the LS algorithm, it has to be implemented by the DFT matrix ${\mathbf{W}}^{\rm DFT}$ with a $Q=M$ pilot length. As for the OMP and VAMP algorithms, we conduct numerical experiments in advance to evaluate their performance under matrices, ${\mathbf{W}}^{\rm Rand}$, ${\mathbf{W}}^{{\rm Top}Q}$, ${\mathbf{W}}^{\rm MM}$ and ${\mathbf{W}}^{\rm IF}$, respectively. It is shown that the combinations of “VAMP + ${\mathbf{W}}^{\rm Rand}$” and “OMP + ${\mathbf{W}}^{{\rm Top}Q}$” exhibit the best channel estimation performance [^2] This is because CS algorithms expect to make the column coherence of observation matrix as small as possible. As the designed matrix ${\mathbf{W}}^{{\rm IF}}$ might select the same eigenvectors in different time slots, it is not suitable for OMP and VAMP algorithms.. Thereby, the following simulations use the random matrix ${\mathbf{W}}^{\rm Rand}$ for VAMP and the top-$Q$ matrix ${\mathbf{W}}^{{\rm Top}Q}$ for OMP. The MMSE algorithm in (4) is the estimator adopted in this paper. We will evaluate its performance under matrices, ${\mathbf{W}}^{\rm Rand}$, ${\mathbf{W}}^{{\rm Top}Q}$, ${\mathbf{W}}^{\rm MM}$, ${\mathbf{W}}^{\rm IF}$, and ${\mathbf{W}}^{\rm WF}$, respectively, to show the superiority of the designed observation matrices.
VI-B Performance Evaluation Under Perfect Kernel
Figure 4: The effect of SNR on NMSE performance under perfect kernel ${\bm{\Sigma}}_{\bf h}$.
In this subsection, we use the perfect kernel, ${\bm{\Sigma}}_{{\mathbf{h}}}$, to trigger the water-filling, ice-filling, and MM algorithms, and then perform the Bayesian regression. The NMSE performance as a function of SNR is evaluated in Fig. 4. The SNR ranges from $-10{\>}{\text{dB}}\sim 10{\>}{\text{dB}}$. Thanks to the careful design of observation matrix, our schemes “MMSE + ${\mathbf{W}}^{\rm MM}$” and “MMSE + ${\mathbf{W}}^{\rm IF}$” remarkably outperform existing benchmarks. For example, a 10 dB NMSE gap between “MMSE + ${\mathbf{W}}^{\rm IF}$” and “VAMP + ${\mathbf{W}}^{\rm Rand}$ is visible. Moreover, the ice-filling matrix ${\mathbf{W}}^{\rm IF}$ can reduce the NMSE of MMSE algorithms by 5 dB compared to the top-$Q$ matrix ${\mathbf{W}}^{{\rm Top}Q}$. Notably, the NMSE performance of MMSE attained by the ice-filling matrix tightly aligns with that of the ideal water-filling matrix, given that the ice-filling intrinsically allocates $Q$ quantized units of power to the $Q$-timeslot pilots. Another finding of interest is that the NMSE gap between “MMSE + ${\mathbf{W}}^{\rm IF}$” and “MMSE + ${\mathbf{W}}^{\rm MM}$” is no higher than 0.5 dB, demonstrating the near-optimality of the proposed MM algorithm in phase-only-controllable design.
Figure 5: The NMSE as a function of pilot length. The pilot length allocated to the LS method is fixed to 128, whereas the pilot length $Q$ for other schemes rises from 8 to 68.
Figure 6: NMSE versus the ratio of antenna spacing and wavelength.
Next, we investigate the impact of pilot length, $Q$, on the NMSE performance in Fig. 5. The pilot length allocated to the LS method is fixed to 128, whereas the pilot length for the other algorithms ranges from 8 to 68. The SNR is -5 dB. It is evident from Fig. 5 that the proposed algorithms consume significantly lower pilot overhead than the other benchmarks to achieve the same NMSE level. For instance, to obtain a $-5\,\text{dB}$ NMSE, the pilot length required for the VAMP algorithm is larger than 30. In contrast, 8 pilots are sufficient for the proposed ice-filling and MM algorithms to attain a NMSE lower than - 10 dB. Additionally, it is notable that the NMSE performance of “MMSE + ${\mathbf{W}}^{{\rm Top}Q}$ no longer decreases when $Q>12$. This is because the top-$Q$ matrix begins to allocate pilots to invalid eigenvectors with very small eigenvalues when $Q>12$, while the proposed ice-filling and MM algorithms continue to select valid eigenvectors from the $1^{\rm st}$ to $12^{\rm th}$ eigenvalues. As a result, the performance gain attained by our algorithms continuously improve as the pilot length increases.
In Fig. 6, we demonstrate the superiority of a DAS in precise channel estimation. The curves of the achieved NMSE performance versus the ratio of antenna spacing and wavelength, i.e., $\frac{d}{\lambda}$, are plotted. The SNR is 0 dB. We can observe that, as the normalized antenna spacing $\frac{d}{\lambda}$ decreases from $\frac{3}{2}$ to $\frac{1}{16}$, the NMSEs for all channel estimators excluding LS method declines rapidly. For example, the NMSE achieved by the proposed ice-filling algorithm is decreased by about 15 dB when the antenna spacing ranges from $\lambda/2$ to $\lambda/8$. This phenomenon is attributed to the fact that, a smaller antenna spacing leads to stronger spatial correlations of channels over different antennas, which results in a more informative kernel ${\bm{\Sigma}}_{{\mathbf{h}}}$ for more precise channel reconstruction. Since the underpinning spatial information of wireless channels is embedded in the structured kernel, these Bayesian regression based channel estimators are enhanced under DASs.
VI-C Performance Evaluation Under Imperfect Kernel
Figure 7: The effect of kernel estimation error on NMSE performance under the imperfect kernel $\hat{\bm{\Sigma}}_{\bf h}$.
In some scenarios where the perfect prior kernel $\bm{\Sigma}_{\bf h}$ cannot be obtained, the statistical kernel and artificial kernels defined in Section IV-E can be utilized to replace the perfect kernel $\bm{\Sigma}_{\bf h}$ in Algorithms 1$\sim$2. We first consider the statistical kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}={\bf{\Sigma}}_{{\bf h}}+\sigma^{2}_{{\mathbf{h}}}{\mathbf{I}}_{M}$. The NMSE as a function of the kernel estimation error $\sigma^{2}_{{\mathbf{h}}}$ is shown in Fig. 7. One can observe that all MMSE methods suffer from visible performance loss due to the imperfect kernel input. Fortunately, the proposed methods still hold the superiority among all schemes. In particular, the estimation accuracy of the proposed methods is almost not influenced when $\sigma^{2}_{{\mathbf{h}}}<-15\,{\text{dB}}$, which shows their high robustness to the kernel error.
Figure 8: The effect of SNR on NMSE performance under the perfect kernel ${\bm{\Sigma}}_{\bf h}$, the statistical kernel $\hat{{\bm{\Sigma}}}_{{\mathbf{h}}}$, the Bessel kernel ${\bm{\Sigma}}_{\rm bes}$, and the exponential kernel ${\bm{\Sigma}}_{\rm exp}$.
We further evaluate the achievable NMSE performance under all considered kernels, including the perfect kernel ${\bm{\Sigma}}_{{\mathbf{h}}}$, the statistical kernel $\hat{{\bm{\Sigma}}}_{{\mathbf{h}}}$, and the artificial kernels $\bm{\Sigma}_{\rm exp}$ and ${\bm{\Sigma}}_{\rm bes}$. The estimation error of $\hat{{\bm{\Sigma}}}_{{\mathbf{h}}}$ is set as ${\mathbf{\sigma}}_{{\mathbf{h}}}^{2}=-12\>{\rm dB}$. The hyper-parameters, $\eta_{1}$ and $\eta_{2}$, of $\bm{\Sigma}_{\rm exp}$ and ${\bm{\Sigma}}_{\rm bes}$ are determined by the binary search through several numerical experiments, which are set to $\eta_{1}=0.56$ and $\eta_{2}=0.85$ in the considered scenario. In this way, we obtain the NMSE versus the SNR and that versus the pilot length $Q$ in Fig. 8 and Fig. 9, respectively. For clarity, we only plot the curves of MMSE algorithms. From these two figures one find that, for a given kernel, the curves of “ideal water-filling”, “proposed ice-filling”, and “proposed MM” are very close, which indicates that these three schemes have similar robustness to imperfect kernels.
[…]
Figure 9: The effect of pilot length on NMSE performance under the perfect kernel ${\bm{\Sigma}}_{\bf h}$, the Bessel kernel ${\bm{\Sigma}}_{\rm bes}$, and the exponential kernel ${\bm{\Sigma}}_{\rm exp}$.
VII Conclusions
This paper incorporated the design of observation matrix into Bayesian channel estimation in DASs. The formulated MIM-based observation matrix design was shown to be a time-domain duality of classic MIMO precoding, which was ideally addressed by the water-filling principle. Targeting practical DAS realizations, we proposed a novel ice-filling and a MM enabled algorithms to design amplitude-and-phase controllable and phase-only controllable observation matrices, respectively. In particular, the ice-filling algorithm was proved to be a discrete approximation of water-filling, with the relative quantization error decaying rapidly. Comprehensive analyses and numerical results validated the near-optimality of the proposed designs.
This work establishes a new understanding of channel estimation from the view of information theory. Several potential directions for extending our work are summarized as follows. In current communication systems, beam selection techniques, including beam training and beam tracking, are widely adopted CSI acquisition approaches. Designing new beam selection strategies using the idea of MIM will be interesting. Besides, the extension to more general systems, such as multi-user MIMO and wideband communications, and theoretically analyzing their channel estimation accuracy also deserve in-depth study.
Appendix A Proof of (9)
The posterior kernel $\bm{\Sigma}_{t}$ is expressed as
${{\bf{\Sigma}}_{t}}={{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t}}{\left({{{\bf{W}}_{t}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t}}}}}+{\sigma^{2}}{{\bf{I}}_{t}}}}\right)^{-1}}{{{{\bf{W}}_{t}^{H}}}}{{\bf{\Sigma}}_{\bf{h}}}.$ (45)
The key to the proof of (9) lies in calculating the inverse of ${{{\bf{W}}_{t}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t}}}}}+{\sigma^{2}}{{\bf{I}}_{t}}}}$ in (45). For ease of notation, we denote ${\bf{K}}_{t}={{{\bf{W}}_{t}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t}}}}}+{\sigma^{2}}{{\bf{I}}_{t}}}}$ and ${\bf{W}}_{t}=[{{\bf{W}}_{t-1}},{\bf{w}}_{t}]$. Then ${\bf{K}}_{t}$ can be expressed in a block matrix form:
${\bf{K}}_{t}=$ $\left[\begin{array}[]{cc}{{{\bf{W}}_{t-1}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}{{{{{\bf{W}}_{t-1}}}}}+{\sigma^{2}}{{\bf{I}}_{t-1}}}}&{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}\\ {\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}&{\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}+\sigma^{2}\end{array}\right]$ (48)
$=$ $\left[\begin{array}[]{cc}{\bf{K}}_{t-1}&{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}\\ {\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}&{\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}+\sigma^{2}\end{array}\right].$ (51)
We can harness the block-matrix inversion in (56)
$\left[\begin{array}[]{cc}{\bf A}&{\bf B}\\ {\bf C}&{\bf D}\end{array}\right]^{-1}=\left[\begin{array}[]{cc}{\bf A}^{-1}+{\bf A}^{-1}{\bf B}({\bf D}-{\bf C}{\bf A}^{-1}{\bf B})^{-1}{\bf C}{\bf A}^{-1}&-{\bf A}^{-1}{\bf B}({\bf D}-{\bf C}{\bf A}^{-1}{\bf B})^{-1}\\ -({\bf D}-{\bf C}{\bf A}^{-1}{\bf B})^{-1}{\bf C}{\bf A}^{-1}&({\bf D}-{\bf C}{\bf A}^{-1}{\bf B})^{-1}\end{array}\right]$ (56)
to derive the inverse of ${\bf{K}}_{t}$ [53]. By mapping the matrices ${\mathbf{A}}$, ${\mathbf{B}}$, ${\mathbf{C}}$, and ${\mathbf{D}}$ in (48) to (56), the value of ${\bf D}-{\bf C}{\bf A}^{-1}{\bf B}$ is expressed as
${\bf D}-{\bf C}{\bf A}^{-1}{\bf B}$
$={\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}+\sigma^{2}-{\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}$
=$\overset{(a)}{=}{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2},$ (57)
where $(a)$ comes from the definition of ${\bf\Sigma}_{t-1}$, i.e., ${\bf\Sigma}_{t-1}={{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}$. Then, by substituting (56) and (A) into (48), the inverse of ${\bf{K}}_{t}$ can be written as (60).
${\bf{K}}_{t}^{-1}=\left[\begin{array}[]{cc}{\bf{K}}_{t-1}^{-1}+\dfrac{{\bf{K}}_{t-1}^{-1}{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}{\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}&-\dfrac{{\bf{K}}_{t-1}^{-1}{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}\vspace*{0.3cm}\\ -\dfrac{{\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}&\dfrac{1}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}\end{array}\right]$ (60)
Taking ${\bf{K}}_{t}^{-1}$ back to the posterior kernel ${\bf\Sigma}_{t}$ in (45) results in
${\bf\Sigma}_{t}=$ ${{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t}}{\bf{K}}_{t}^{-1}{{{{\bf{W}}_{t}^{H}}}}{{\bf{\Sigma}}_{\bf{h}}}$
$=$ ${{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}{{{{\bf{W}}_{t-1}^{H}}}}{{\bf{\Sigma}}_{\bf{h}}}+$
$\frac{1}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}({\bf E}+{\bf F}+{\bf G}+{\bf H})$
$=$ ${\bf\Sigma}_{t-1}+\frac{1}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}({\bf E}+{\bf F}+{\bf G}+{\bf H}),$ (61)
where
${\bf E}$ $=-{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}{{{{\bf{W}}_{t-1}^{H}}}}{{\bf{\Sigma}}_{\bf{h}}}{\bf w}_{t}{\bf w}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}{{{{\bf{W}}_{t-1}^{H}}}}{{\bf{\Sigma}}_{\bf{h}}}$
$=-({{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{t-1}}){\bf w}_{t}{\bf w}_{t}^{H}({{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{t-1}}),$
${\bf F}$ $={{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}{{{{\bf{W}}_{t-1}^{H}}}}{{\bf{\Sigma}}_{\bf{h}}}{\bf w}_{t}{\bf w}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}$
$=({{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{t-1}}){\bf w}_{t}{\bf w}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}},$
${\bf G}$ $={{\bf{\Sigma}}_{\bf{h}}}{\bf w}_{t}{\bf w}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}{{{{\bf{W}}_{t-1}^{H}}}}{{\bf{\Sigma}}_{\bf{h}}}$
$={{\bf{\Sigma}}_{\bf{h}}}{\bf w}_{t}{\bf w}_{t}^{H}({{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{t-1}}),$
${\bf H}$ $=-{{\bf{\Sigma}}_{\bf{h}}}{\bf w}_{t}{\bf w}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}.$
It is straightforward to verify that the sum of ${\bf E}$, ${\bf F}$, ${\bf G}$, and ${\bf H}$ is exactly $-{{\bf{\Sigma}}_{t-1}}{\bf w}_{t}{\bf w}_{t}^{H}{{\bf{\Sigma}}_{t-1}}$. By substituting ${\bf E}$, ${\bf F}$, ${\bf G}$, and ${\bf H}$ into (A), we have
${\bf{\Sigma}}_{t}={\bf{\Sigma}}_{t-1}-\frac{{\bf{\Sigma}}_{t-1}{\bf{w}}_{t}{\bf{w}}_{t}^{H}{\bf{\Sigma}}_{t-1}}{{\bf{w}}_{t}^{H}{\bf{\Sigma}}_{t-1}{\bf{w}}_{t}+\sigma^{2}},$ (62)
which completes the proof.
Appendix B Proof of Theorem 2
Theorem 2 can be proved by the contradiction method. Suppose there exists an index $k$ such that $n_{k}\geq p_{k}+1$. Due to the constraint that $\sum_{k=1}^{K}n_{k}=\sum_{k=1}^{K}p_{k}=Q$, there must exists $k^{\prime}$ such that $k^{\prime}\neq k$ and $n_{k^{\prime}}<p_{k^{\prime}}$, otherwise $\sum_{k=1}^{K}n_{k}$ will be larger than $\sum_{k=1}^{K}p_{k}$. Note that $p_{k^{\prime}}$ is non-zero since $p_{k^{\prime}}>n_{k^{\prime}}\geq 0$. In this context, the water level can be expressed as $\beta=p_{k^{\prime}}+\frac{\sigma^{2}}{\lambda_{k^{\prime}}}$ and the $k^{\prime}$-th ice level $\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{Q}}$ should be smaller than $\beta$ because
$\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{Q}}=n_{k^{\prime}}+\frac{\sigma^{2}}{\lambda_{k^{\prime}}}<p_{k^{\prime}}+\frac{\sigma^{2}}{\lambda_{k^{\prime}}}=\beta.$ (63)
Then, consider the $k$-th ice level $\frac{\sigma^{2}}{\lambda_{k}^{Q}}$. Since $n_{k}\geq p_{k}+1$, we have the following inequality:
$\frac{\sigma^{2}}{\lambda_{k}^{Q}}$ $=n_{k}+\frac{\sigma^{2}}{\lambda_{k}}\geq p_{k}+\frac{\sigma^{2}}{\lambda_{k}}+1=\left(\beta-\frac{\sigma^{2}}{\lambda_{k}}\right)^{+}+\frac{\sigma^{2}}{\lambda_{k}}+1$
$\overset{(a)}{\geq}\beta+1,$ (64)
where the inequality (a) holds because $\left(\beta-\frac{\sigma^{2}}{\lambda_{k}}\right)^{+}\geq\beta-\frac{\sigma^{2}}{\lambda_{k}}$. Combining equations (63) and (B), we arrive at
$\frac{\sigma^{2}}{\lambda_{k}^{Q}}>\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{Q}}+1.$ (65)
We introduce the following Lemma 2.1 to evaluate the inequality (65).
Lemma 2.1: For the $k$-th eigenvalue $\lambda_{k}^{t}$ obtained at the $t$-th timeslot, if $n_{k}^{t}>0$, we have the inequality
$\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{t}}+1\geq\frac{\sigma^{2}}{\lambda_{k}^{t}},$ (66)
hold for all $k^{\prime}\in\{1,2,\cdots,K\}$.
Proof:
Since $n_{k}^{t}>0$, the $k$-th eigenvector is selected by the ice-filling algorithm at least once. Let $t^{\prime}<t$ denote the latest timeslot before $t$ when the $k$-th eigenvector is selected for pilot transmission such that $n_{k}^{t^{\prime}}=n_{k}^{t}-1$ and $n_{k}^{t^{\prime}+1}=n_{k}^{t}$.
Suppose there exists $k^{\prime}$ such that $\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{t}}<\frac{\sigma^{2}}{\lambda_{k}^{t}}-1$, then we have
$\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{t^{\prime}}}\overset{(b)}{\leq}\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{t}}\overset{(c)}{<}\frac{\sigma^{2}}{\lambda_{k}}+n_{k}^{t}-1=\frac{\sigma^{2}}{\lambda_{k}}+n_{k}^{t^{\prime}}=\frac{\sigma^{2}}{\lambda_{k}^{t^{\prime}}},$ (67)
where (b) holds because the ice-level $\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{t}}$ is non-decreasing w.r.t the timeslot $t$, and (c) holds given that $\frac{\sigma^{2}}{\lambda_{k}}+n_{k}^{t}=\frac{\sigma^{2}}{\lambda_{k}^{t}}$. The inequality $\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{t^{\prime}}}<\frac{\sigma^{2}}{\lambda_{k}^{t^{\prime}}}$ in (67) contradicts the eigenvector selection principle of the ice-filling algorithm that $\lambda_{k}^{t^{\prime}}$ should be the largest over the eigenvalue set $\{\lambda_{1}^{t^{\prime}},\lambda_{2}^{t^{\prime}},\cdots,\lambda_{K}^{t^{\prime}}\}$. Therefore, $\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{t}}$ should be no less than $\frac{\sigma^{2}}{\lambda_{k}^{t}}-1$ for all $k^{\prime}$, which completes the proof. ∎
Comparing (65) and (66), it is clear that (65) contradicts Lemma 2.1 given that $n_{k}\geq p_{k}+1>0$. Therefore, $n_{k}$ should be smaller than $p_{k}+1$.
We adopt the same method to prove that $n_{k}>p_{k}-1$. Specifically, suppose there is $k$ such that $n_{k}\leq p_{k}-1$. Then there must exist $k^{\prime}\neq k$ such that $n_{k^{\prime}}>p_{k^{\prime}}$. From the inequality $n_{k}\leq p_{k}-1$, we know that the ice level $\frac{\sigma^{2}}{\lambda_{k}^{Q}}$ satisfies the inequality
$\frac{\sigma^{2}}{\lambda_{k}^{Q}}=n_{k}+\frac{\sigma^{2}}{\lambda_{k}}\leq p_{k}+\frac{\sigma^{2}}{\lambda_{k}}-1\overset{(b)}{=}\beta-1,$ (68)
where (b) arises because $p_{k}$ is greater than 0. On the other hand, from the inequality $n_{k^{\prime}}>p_{k^{\prime}}$, we get
$\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{Q}}=n_{k^{\prime}}+\frac{\sigma^{2}}{\lambda_{k^{\prime}}}>p_{k^{\prime}}+\frac{\sigma^{2}}{\lambda_{k^{\prime}}}\geq\beta.$ (69)
Combining (68) and (69), we get
$\frac{\sigma^{2}}{\lambda_{k^{\prime}}^{Q}}\geq\frac{\sigma^{2}}{\lambda_{k}^{Q}}+1,$ (70)
which contradicts Lemma 2.1 because $n_{k^{\prime}}>p_{k^{\prime}}\geq 0$. As a result, $n_{k}$ should be larger than $p_{k}-1$, and thereby we get
$|n_{k}-p_{k}|<1,$ (71)
which completes the proof.
Appendix C Proof of Lemma 1
Substituting ${\mathbf{W}}={\mathbf{U}}_{K}{\mathbf{P}}$ and ${\bf\Sigma_{h}}={\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}$ into the definition of ${{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}$, we get
${{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}={{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}}{\left({\bf W}^{H}{\bf{\Sigma}}_{\bf{h}}{\bf W}+\sigma^{2}{\bf I}_{Q}\right)^{-1}}{{{{\bf{W}}^{H}}}}{{\bf{\Sigma}}_{\bf{h}}}$
$=$ ${\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}-{\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}{\mathbf{U}}_{K}{\mathbf{P}}({\mathbf{P}}^{H}{\mathbf{U}}_{K}^{H}{\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}{\mathbf{U}}_{K}{\mathbf{P}}$
$+\sigma^{2}{\mathbf{I}}_{Q})^{-1}{\mathbf{P}}^{H}{\mathbf{U}}_{K}^{H}{\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}$
$=$ ${\mathbf{U}}_{K}\underbrace{\left({\bf\Lambda}_{K}-{\bf\Lambda}_{K}{\mathbf{P}}\left({\mathbf{P}}^{H}{\bf\Lambda}_{K}{\mathbf{P}}+\sigma^{2}{\mathbf{I}}_{Q}\right)^{-1}{\mathbf{P}}^{H}{\bf\Lambda}_{K}\right)}_{{\bf\Lambda}_{{\mathbf{h}}|{\mathbf{y}}}}{\mathbf{U}}_{K}^{H}.$ (72)
Note that matrices ${\bf\Lambda}_{K}$ and ${\mathbf{P}}$ are both diagonal. Hence, it can be directly proved that
${\bf\Lambda}_{{\mathbf{h}}|{\mathbf{y}}}={\rm diag}\left(\frac{\lambda_{1}\sigma^{2}}{p_{1}\lambda_{1}+\sigma^{2}},\cdots,\frac{\lambda_{K}\sigma^{2}}{p_{K}\lambda_{K}+\sigma^{2}}\right).$ (73)
Therefore, given that ${\mathbf{U}}_{K}^{H}{\mathbf{U}}_{K}={\mathbf{I}}_{K}$, the MSE achieved by water-filling can be derived as
$\mathsf{Tr}\left({{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}\right)=\mathsf{Tr}\left({\bf\Lambda}_{{\mathbf{h}}|{\mathbf{y}}}\right)=\sum_{k=1}^{K}\frac{\lambda_{k}\sigma^{2}}{p_{k}\lambda_{k}+\sigma^{2}}.$ (74)
Particularly, as $Q\rightarrow+\infty$, the water-filling principle asymptotically allocates equal power $Q/K$ to all eigenvectors [54], rendering the estimation error decays at a rate of $\mathcal{O}\left(\sum_{k=1}^{K}\frac{K\sigma^{2}}{Q}\right)=\mathcal{O}\left(K^{2}Q^{-1}\right)$. This completes the proof.
Appendix D Proof of Lemma 2
We first express the observation matrix designed by the ice-filling algorithm as ${\mathbf{W}}={\mathbf{U}}_{K}{\mathbf{S}}$. Here, ${\mathbf{S}}\in\{0,1\}^{K\times Q}$ represents a switch matrix, each column of which contains one non-zero entry, i.e., ${\mathbf{S}}(k,q)\in\{0,1\}$ and $\sum_{k=1}^{K}{\mathbf{S}}(k,q)=1$. In particular, ${\mathbf{S}}(k,q)=1$ implies that the $k$-th eigenvector, ${\mathbf{u}}_{k}$, is assigned to the $q$-th observation vector. Based on this definition, we can prove that
${\mathbf{S}}{\mathbf{S}}^{H}={\mathbf{N}}={\rm diag}(n_{1},n_{2},\cdots,n_{K}),$ (75)
given that $n_{k}$ ($k\in\{1,2,\cdots,K\}$) stands for the number of times that the eigenvector ${\mathbf{u}}_{k}$ is assigned to ${\mathbf{W}}$ by Algorithm 1. Then, by substituting ${\mathbf{W}}={\mathbf{U}}_{K}{\mathbf{S}}$ and ${\bf\Sigma_{h}}={\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}$ into (45), we can express ${{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}$ as
${{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}={\mathbf{U}}_{K}{\bf\Lambda}_{{\mathbf{h}}|{\mathbf{y}}}{\mathbf{U}}_{K}^{H},$ (76)
where ${\bf\Lambda}_{{\mathbf{h}}|{\mathbf{y}}}={\bf\Lambda}_{K}-{\bf\Lambda}_{K}{\mathbf{S}}\left({\mathbf{S}}^{H}{\bf\Lambda}_{K}{\mathbf{S}}+\sigma^{2}{\mathbf{I}}_{Q}\right)^{-1}{\mathbf{S}}^{H}{\bf\Lambda}_{K}$. Moreover, the Sherman-Morrison-Woodbury formula [53]
$({\mathbf{A}}+{\mathbf{X}}{\mathbf{R}}{\mathbf{Y}})^{-1}={\mathbf{A}}^{-1}-{\mathbf{A}}^{-1}{\mathbf{X}}({\mathbf{R}}^{-1}+{\mathbf{Y}}{\mathbf{A}}^{-1}{\mathbf{X}})^{-1}{\mathbf{Y}}{\mathbf{A}}^{-1}$ (77)
allows us to rewrite ${\bf\Lambda}_{{\mathbf{h}}|{\mathbf{y}}}$ as
${\bf\Lambda}_{{\mathbf{h}}|{\mathbf{y}}}={\bf\Lambda}_{K}-\frac{1}{\sigma^{2}}{\bf\Lambda}_{K}{\mathbf{S}}{\mathbf{S}}^{H}{\bf\Lambda}_{K}$
$~{}~{}~{}~{}~{}~{}+\frac{1}{\sigma^{4}}{\bf\Lambda}_{K}{\mathbf{S}}{\mathbf{S}}^{H}\left({\bf\Lambda}_{K}^{-1}+\frac{1}{\sigma^{2}}{\mathbf{S}}{\mathbf{S}}^{H}\right)^{-1}{\mathbf{S}}{\mathbf{S}}^{H}{\bf\Lambda}_{K}$
=$\overset{(a)}{=}{\bf\Lambda}_{K}\!-\!\frac{1}{\sigma^{2}}{\bf\Lambda}_{K}{\mathbf{N}}{\bf\Lambda}_{K}\!+\!\frac{1}{\sigma^{4}}{\bf\Lambda}_{K}{\mathbf{N}}\left({\bf\Lambda}_{K}^{-1}\!+\!\frac{1}{\sigma^{2}}{\mathbf{N}}\right)^{\!\!-1}\!\!{\mathbf{N}}{\bf\Lambda}_{K},$
=$\overset{(b)}{=}{\rm diag}\left(\frac{\lambda_{1}\sigma^{2}}{n_{1}\lambda_{1}+\sigma^{2}},\cdots,\frac{\lambda_{K}\sigma^{2}}{n_{K}\lambda_{K}+\sigma^{2}}\right),$ (78)
where $(a)$ holds since ${\mathbf{S}}{\mathbf{S}}^{H}={\mathbf{N}}$, and the equation $(b)$ comes directly from the fact the matrices ${\mathbf{\Lambda}}_{K}$ and ${\mathbf{N}}$ are both diagonal. In conclusion, the MSE achieved by the ice-filling algorithm is $\mathsf{Tr}({\bf\Lambda}_{{\mathbf{h}}|{\mathbf{y}}})=\sum_{k=1}^{K}\frac{\lambda_{k}\sigma^{2}}{n_{k}\lambda_{k}+\sigma^{2}}$. Moreover, since $|n_{k}-p_{k}|<1$, the asymptotic estimation error is naturally $\mathcal{O}(K^{2}Q^{-1})$ as $Q\rightarrow+\infty$, which completes the proof.
Appendix E Proof of Theorem 3
We use the Theorem 2 to upper-bound the asymptotic MSE difference as $Q\rightarrow+\infty$, i.e.,
$|\delta_{\rm wf}-\delta_{\rm if}|$ $=\left|\sum_{k=1}^{K}\frac{\lambda_{k}\sigma^{2}}{p_{k}\lambda_{k}+\sigma^{2}}-\sum_{k=1}^{K}\frac{\lambda_{k}\sigma^{2}}{n_{k}\lambda_{k}+\sigma^{2}}\right|$
$\leq\sum_{k=1}^{K}\left|\frac{\lambda_{k}\sigma^{2}}{p_{k}\lambda_{k}+\sigma^{2}}-\frac{\lambda_{k}\sigma^{2}}{n_{k}\lambda_{k}+\sigma^{2}}\right|$
$=\sum_{k=1}^{K}\left|\frac{\lambda_{k}^{2}\sigma^{2}(p_{k}-n_{k})}{(p_{k}\lambda_{k}+\sigma^{2})(n_{k}\lambda_{k}+\sigma^{2})}\right|$
<$\overset{(a)}{<}\sum_{k=1}^{K}\left|\frac{\lambda_{k}^{2}\sigma^{2}}{(p_{k}\lambda_{k}+\sigma^{2})((p_{k}-1)\lambda_{k}+\sigma^{2})}\right|,$ (79)
where (a) holds due to the facts that $|n_{k}-p_{k}|<1$ and $n_{k}>p_{k}-1>0$ when $Q\rightarrow+\infty$. Moreover, the water-filling principle tends to allocate equal power to all channels, i.e., $p_{k}\rightarrow Q/K$ for large $Q$ [54]. Thereafter, the asymptotic upper-bound of the MSE difference becomes
$|\delta_{\rm wf}-\delta_{\rm if}|<\sum_{k=1}^{K}\frac{K^{2}\sigma^{2}}{Q^{2}}=\mathcal{O}(K^{3}Q^{-2}).$ (80)
Appendix F Proof of Lemma 3
Let ${\bf\Sigma_{h}}={\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}$ denote the eigenvalue decomposition of the rank-$K$ kernel ${{\bf{\Sigma}}_{\bf{h}}}$, wherein ${\mathbf{U}}_{K}\in{\mathbb{C}}^{M\times K}$ and ${\mathbf{\Lambda}}_{K}\in{\mathbb{C}}^{M\times M}$. According to the extension of the Sherman-Morrison-Woodbury formula in (77), the MSE achieved by the proposed algorithms can be rewritten as
$\mathsf{E}\left(\|{\bm{\mu}_{{\mathbf{h}}|{\mathbf{y}}}}-{\mathbf{h}}\|^{2}_{2}\right)=\mathsf{Tr}\left({{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}\right)$
$=$ $\mathsf{Tr}\left({{{\bf{U}}_{K}}{\left({{\bf{\Lambda}}_{K}^{-1}+{1\over{{\sigma^{2}}}}{\bf{U}}_{K}^{H}{\bf{W}}{{\bf{W}}^{H}}{{\bf{U}}_{K}}}\right)^{-1}}{\bf{U}}_{K}^{H}}\right).$ (81)
For the randomly generated observation matrix $\bf W$ in both the amplitude-and-phase controllable and phase-only controllable cases, following the asymptotic orthogonality principle, we have ${\bf{W}}{{\bf{W}}^{H}}\approx{Q\over M}{{\bf{I}}_{M}}$ when $Q$ is sufficiently large. In this case, the posterior kernel can be approximated by
${{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}\approx{{\bf{U}}_{K}}{\left({{\bf{\Lambda}}_{K}^{-1}+{1\over{{\sigma^{2}}}}{Q\over M}{\bf{U}}_{K}^{H}{{\bf{U}}_{K}}}\right)^{-1}}{\bf{U}}_{K}^{H}=$
${{\bf{U}}_{K}}{\rm{diag}}\left({{1\over{{1\over{{\lambda_{1}}}}+{Q\over{M{\sigma^{2}}}}}},{1\over{{1\over{{\lambda_{2}}}}+{Q\over{M{\sigma^{2}}}}}},\cdots,{1\over{{1\over{{\lambda_{K}}}}+{Q\over{M{\sigma^{2}}}}}}}\right){\bf{U}}_{K}^{H}.$ (82)
Therefore, by substituting (F) into (F), the MSE $\delta_{\rm rnd}$ is given by
$\delta_{\rm rnd}\approx\sum\limits_{k=1}^{K}{{1\over{{1\over{{\lambda_{k}}}}+{Q\over{M{\sigma^{2}}}}}}}=\sum\limits_{k=1}^{K}{{{{\lambda_{k}}{\sigma^{2}}}\over{{Q\over M}{\lambda_{k}}+{\sigma^{2}}}}},$ (83)
which completes the proof.
Appendix G Proof of Lemma 4
Given a prior kernel $\bf\Sigma$ as the input of the proposed estimator, the squared error of channel estimation can be derived as
${{{\left\|{{{\bm{\mu}}}_{{\mathbf{h}}|{\mathbf{y}}}-{\bf{h}}}\right\|}^{2}_{2}}}={\left\|{{\bf{\Sigma W}}{{\left({{{\bf{W}}^{H}}{\bf{\Sigma W}}+{\sigma^{2}}{{\bf{I}}_{Q}}}\right)}^{-1}}{\bf{y}}-{\bf{h}}}\right\|^{2}_{2}}$
=$\stackrel{{\scriptstyle(a)}}{{=}}{\Big{\|}}\left({{\bf{\Sigma W}}{{\left({{{\bf{W}}^{H}}{\bf{\Sigma W}}+{\sigma^{2}}{\bf I}_{Q}}\right)}^{-1}}{{\bf{W}}^{H}}-{{\bf{I}}_{M}}}\right){{\bf{h}}}$
$~{}~{}~{}~{}~{}~{}~{}~{}+{\bf{\Sigma W}}{\left({{{\bf{W}}^{H}}{\bf{\Sigma W}}+{\sigma^{2}}{\bf I}_{Q}}\right)^{-1}}{\bf{z}}{\Big{\|}}^{2}_{2}$
=$\stackrel{{\scriptstyle(b)}}{{=}}{\left\|\left({{{\bf{\Pi}}^{H}}{{\bf{W}}^{H}}-{{\bf{I}}_{M}}}\right){\bf{h}}+{{\bf{\Pi}}^{H}}{\bf{z}}\right\|^{2}_{2}},$ (84)
where $(a)$ holds since ${\bf y}={\bf W}^{H}{\bf h}+{\bf z}$ and $(b)$ holds by defining ${\bf{\Pi}}={\left({{{\bf{W}}^{H}}{\bf{\Sigma W}}+{\sigma^{2}}{\bf I}_{Q}}\right)^{-1}}{{\bf{W}}^{H}}{\bf{\Sigma}}$, which can be viewed as a matrix function with respect to the artificial kernel $\bf\Sigma$. Next, recalling properties ${\bf z}\sim\mathcal{C}\mathcal{N}\!\left({\bf 0}_{Q},\sigma^{2}{\bf I}_{Q}\right)$ and ${\mathsf{E}}\left({\bf h}{\bf h}^{\rm H}\right)={\bf{\Sigma}}_{{\bf h}}$, the MSE ${\hat{\delta}}$ under imperfect kernel can be derived as
${\hat{\delta}}=\mathsf{Tr}\left(\left({{{\bf{\Pi}}^{H}}{{\bf{W}}^{H}}-{{\bf{I}}_{M}}}\right)\!{{\bf{\Sigma}}_{\bf{h}}}\!\left({{\bf{W\Pi}}-{{\bf{I}}_{M}}}\right)\right)+{\sigma^{2}}\mathsf{Tr}\!\left({{\bf{\Pi}}^{H}}{\bf{\Pi}}\right),$ (85)
where the equality holds since ${\left\|{\bf{x}}\right\|^{2}_{2}}={\mathsf{Tr}}\left({{\bf{x}}{{\bf{x}}^{\rm H}}}\right)$ for any vector $\bf x$, which completes the proof.
Appendix H Proof of Lemma 5
Comparing (5) and (34), we only need to prove that ${\bf{W\Pi}}={\bf{\Omega}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}$ and ${{\bf{\Pi}}^{H}}{\bf{\Pi}}={{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\bf{\Xi}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}$ in the case when ${\bf{\Sigma}}={\bf{\hat{\Sigma}}}_{{\bf h}}$. Due to the positive semi-definition of ${\bf{\Sigma}}_{{\bf h}}$, the inversion of ${\bf{\hat{\Sigma}}}_{{\bf h}}={\bf{\Sigma}}_{{\bf h}}+\sigma^{2}_{\bf h}{\bf I}_{M}$ exists. Thus, by applying the Sherman-Morrison-Woodbury formula in (77) for the term ${\left({{{\bf{W}}^{H}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\bf{W}}+{\sigma^{2}}{{\bf{I}}_{Q}}}\right)^{-1}}$, ${\bf{W\Pi}}$ can be reorganized as
${\bf{W}}{{\bf{\Pi}}}={\bf{W}}{\left({{{\bf{W}}^{H}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\bf{W}}+{\sigma^{2}}{{\bf{I}}_{Q}}}\right)^{-1}}{{\bf{W}}^{H}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}$
$={\bf{W}}\left({{1\over{{\sigma^{2}}}}{{\bf{I}}_{Q}}-{1\over{{\sigma^{4}}}}{{\bf{W}}^{H}}{{\left({{\bf{\hat{\Sigma}}}_{\bf{h}}^{-1}\!+\!{1\over{{\sigma^{2}}}}{\bf{W}}{{\bf{W}}^{H}}}\right)}^{\!-1}}{\bf{W}}}\right){{\bf{W}}^{H}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}$
$=\left({{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{2}}}}\!-\!{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{4}}}}{{\left({{\bf{\hat{\Sigma}}}_{\bf{h}}^{\!-1}\!+\!{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{2}}}}}\right)}^{-1}}{\bf{W}}{{\bf{W}}^{H}}}\right){{{\bf{\hat{\Sigma}}}}_{\bf{h}}},$ (86)
which is equal to ${\bm{\Omega}}{{\bf{\hat{\Sigma}}}}_{\bf{h}}$ and obviously the is a function of ${\bf W}{\bf W}^{H}$. Using the same matrix techniques as (H), we can prove that
${{\bf{\Pi}}^{H}}{\bf{\Pi}}={{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\bf{W}}{\left({{{\bf{W}}^{H}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\bf{W}}+{\sigma^{2}}{{\bf{I}}_{Q}}}\right)^{-2}}{{\bf{W}}^{H}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}$
$=$ ${{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\bf{W}}{\left({{1\over{{\sigma^{2}}}}{{\bf{I}}_{Q}}\!-\!{1\over{{\sigma^{4}}}}{{\bf{W}}^{H}}{{\left({{\bf{\hat{\Sigma}}}_{\bf{h}}^{-1}\!+\!{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{2}}}}}\right)}^{\!\!-1}}{\bf{W}}}\right)^{\!\!2}}{{\bf{W}}^{H}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}$
$=$ ${{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\Bigg{(}}{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{4}}}}-2{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{6}}}}{\left({{\bf{\hat{\Sigma}}}_{\bf{h}}^{-1}+{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{2}}}}}\right)^{-1}}{\bf{W}}{{\bf{W}}^{H}}+$
${{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{8}}}}\left({\left({{\bf{\hat{\Sigma}}}_{\bf{h}}^{-1}+{{{\bf{W}}{{\bf{W}}^{H}}}\over{{\sigma^{2}}}}}\right)^{-1}}{\bf{W}}{{\bf{W}}^{H}}\right)^{2}{\Bigg{)}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}},$ (87)
which is equal to ${{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\bf{\Xi}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}$ and obviously is a function of ${\bf W}{\bf W}^{H}$. This completes the proof.
Appendix I Proof of Lemma 6
Let ${\bf\Sigma_{h}}={\mathbf{U}}{\bf\Lambda}{\mathbf{U}}^{H}$ denote the complete eigenvalue decomposition of the kernel ${{\bf{\Sigma}}_{\bf{h}}}$, wherein ${\mathbf{U}}\in{\mathbb{C}}^{M\times M}$ and ${\mathbf{\Lambda}}=\text{diag}\{\lambda_{1},\cdots,\lambda_{M}\}$ with $\lambda_{m}=0$ for $K<m\leq M$. According to the analyses in Appendices C, D, and F, the term ${\bf W}{\bf W}^{H}$ for the water-filling estimator, ice-filling estimator, and random estimator can be respectively written as
${{\bf{W}}_{{\rm{wf}}}}{\bf{W}}_{{\rm{wf}}}^{H}={\bf{U\hat{P}}}{{{\bf{\hat{P}}}}^{H}}{{\bf{U}}^{H}},$ (88a)
${{\bf{W}}_{{\rm{if}}}}{\bf{W}}_{{\rm{if}}}^{H}={\bf{U\hat{N}}}{{\bf{U}}^{H}},$ (88b)
${{\bf{W}}_{{\rm{rnd}}}}{\bf{W}}_{{\rm{rnd}}}^{H}\approx{Q\over M}{{\bf{I}}_{M}}={\bf{U}}\left({{Q\over M}{{\bf{I}}_{M}}}\right){{\bf{U}}^{H}}.$ (88c)
where ${\hat{\bf P}}={\rm diag}\left(\sqrt{{\hat{p}}_{1}},\cdots,\sqrt{{\hat{p}}_{M}}\right)$ is the power allocated to the $M$ eigenvectors of $\hat{\bf\Sigma}_{\bf h}$ and ${\hat{\bf N}}={\rm diag}\left({{\hat{n}}_{1}},\cdots,{{\hat{n}}_{M}}\right)$ is the pilot-reuse-frequency matrix, as defined in (75). It is clear from (88) that the terms ${\bf W}{\bf W}^{H}$ in the three cases share a general form ${\bf W}{\bf W}^{H}={\bf{U}}{\bm{\Psi}}{\bf{U}}^{H}$ wherein ${\bm{\Psi}}={\rm diag}\left(\psi_{1},\cdots,\psi_{M}\right)$ is a diagonal matrix. Therefore, the proof of Lemma 6 can be obtained by replacing ${\bm{\Psi}}$ with the corresponding terms ${\bf{\hat{P}}}{{{\bf{\hat{P}}}}^{H}}$, ${\bf{\hat{N}}}$, and ${{Q\over M}{{\bf{I}}_{M}}}$,respectively. It allows us focus on the general proof based on the common expression ${\bf{U}}{\bm{\Psi}}{\bf{U}}^{H}$, which is given as follows.
Recalling the useful expression in Lemma 5, by substituting ${\bf W}{\bf W}^{H}={\bf{U}}{\bm{\Psi}}{\bf{U}}^{H}$, ${\bf\Sigma_{h}}={\mathbf{U}}{\bf\Lambda}{\mathbf{U}}^{H}$, and ${\bf{\hat{\Sigma}}}_{{\bf h}}={\bf{\Sigma}}_{{\bf h}}+\sigma^{2}_{\bf h}{\bf I}_{M}$ into $\bm{\Omega}$ in (36) and $\bm{\Xi}$ in (37), we arrive at
${\bm{\Omega}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}={\bf{U}}\left({{{\bf{\Psi}}\over{{\sigma^{2}}}}-{{\bf{\Psi}}\over{{\sigma^{4}}}}{{\left({{{\left({{\bf{\Lambda}}+\sigma_{\bf{h}}^{2}{{\bf{I}}_{M}}}\right)}^{-1}}+{{\bf{\Psi}}\over{{\sigma^{2}}}}}\right)}^{-1}}{\bf{\Psi}}}\right){{\bf\Lambda}{\mathbf{U}}^{H}}$
$={\bf{U}}{\rm{diag}}\!\left({{{{\psi_{1}}\left({{\lambda_{1}}\!+\!\sigma_{\bf{h}}^{2}}\right)}\over{{\sigma^{2}}\!+\!{\psi_{1}}\left({{\lambda_{1}}\!+\!\sigma_{\bf{h}}^{2}}\right)}},\cdots,{{{\psi_{M}}\left({{\lambda_{M}}\!+\!\sigma_{\bf{h}}^{2}}\right)}\over{{\sigma^{2}}\!+\!{\psi_{M}}\left({{\lambda_{M}}\!+\!\sigma_{\bf{h}}^{2}}\right)}}}\right){{\bf{U}}^{H}}$
$={\bf{U}}{\bm{\Lambda}}_{\bm{\Omega}}{{\bf{U}}^{H}},$ (89)
${{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\bf{\Xi}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}={\bf{U}}\left({{\bf{\Lambda}}+\sigma_{\bf{h}}^{2}{{\bf{I}}_{M}}}\right)\times$
${\Bigg{(}}{{\bf{\Psi}}\over{{\sigma^{4}}}}-2{{\bf{\Psi}}\over{{\sigma^{6}}}}{\left({{{\left({{\bf{\Lambda}}+\sigma_{\bf{h}}^{2}{{\bf{I}}_{L}}}\right)}^{-1}}+{{\bf{\Psi}}\over{{\sigma^{2}}}}}\right)^{-1}}{\bf{\Psi}}+$
${{\bf{\Psi}}\over{{\sigma^{8}}}}{\left({{{\left({{{\left({{\bf{\Lambda}}+\sigma_{\bf{h}}^{2}{{\bf{I}}_{L}}}\right)}^{-1}}+{{\bf{\Psi}}\over{{\sigma^{2}}}}}\right)}^{-1}}{\bf{\Psi}}}\right)^{2}}{\Bigg{)}}\left({{\bf{\Lambda}}+\sigma_{\bf{h}}^{2}{{\bf{I}}_{M}}}\right){{\bf{U}}^{H}}=$
${\bf{U}}{\rm{diag}}\!\left({{{{\psi_{1}}{{\left({{\lambda_{1}}\!+\!\sigma_{\bf{h}}^{2}}\right)}^{2}}}\over{{{\left({{\psi_{1}}\left({{\lambda_{1}}\!+\!\sigma_{\bf{h}}^{2}}\right)\!+\!{\sigma^{2}}}\right)}^{2}}}},\cdots,{{{\psi_{M}}{{\left({{\lambda_{M}}\!+\!\sigma_{\bf{h}}^{2}}\right)}^{2}}}\over{{{\left({{\psi_{M}}\left({{\lambda_{M}}\!+\!\sigma_{\bf{h}}^{2}}\right)\!+\!{\sigma^{2}}}\right)}^{2}}}}}\right)\!{{\bf{U}}^{H}}$
$={\bf{U}}{\bm{\Lambda}}_{\bm{\Xi}}{{\bf{U}}^{H}},$ (90)
where ${\bm{\Lambda}}_{\bm{\Omega}}$ and ${\bm{\Lambda}}_{\bm{\Xi}}$ are both $M$-dimensional diagonal matrices. By substituting the above ${\bm{\Omega}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}$, ${{{\bf{\hat{\Sigma}}}}_{\bf{h}}}{\bf{\Xi}}{{{\bf{\hat{\Sigma}}}}_{\bf{h}}}$, and ${\bf\Sigma_{h}}={\mathbf{U}}{\bf\Lambda}{\mathbf{U}}^{H}$ into $\hat{\delta}$ in (5), we obtain
$\hat{\delta}=\mathsf{Tr}\left({\left({{{\bf{\Lambda}}_{\bf{\Omega}}}-{{\bf{I}}_{M}}}\right){\bf{\Lambda}}\left({{{\bf{\Lambda}}_{\bf{\Omega}}}-{{\bf{I}}_{M}}}\right)}\right)+{\sigma^{2}}\mathsf{Tr}\left({{{\bf{\Lambda}}_{\bf{\Xi}}}}\right)$
$=$ ${\sigma^{2}}\sum\limits_{m=1}^{M}{{{{\lambda_{m}}{\sigma^{2}}+{\psi_{m}}{{\left({{\lambda_{m}}+\sigma_{\bf{h}}^{2}}\right)}^{2}}}\over{{{\left({{\psi_{m}}\left({{\lambda_{m}}+\sigma_{\bf{h}}^{2}}\right)+{\sigma^{2}}}\right)}^{2}}}}}$
$=$ ${\sigma^{2}}\sum\limits_{k=1}^{K}{\frac{{{\lambda_{k}}{\sigma^{2}}+{\psi_{k}}{{\left({{\lambda_{k}}+\sigma_{\bf{h}}^{2}}\right)}^{2}}}}{{{{\left({{\psi_{k}}\left({{\lambda_{k}}+\sigma_{\bf{h}}^{2}}\right)+{\sigma^{2}}}\right)}^{2}}}}}+\sum\limits_{m=K+1}^{M}{\frac{{{\psi_{m}}\sigma^{2}\sigma_{\bf{h}}^{4}}}{{{{\left({{\psi_{m}}\sigma_{\bf{h}}^{2}+{\sigma^{2}}}\right)}^{2}}}}},$ (91)
where the second equality holds since $\lambda_{m}=0$ for all $m>K$. According to (88), for the water-filling estimator, ice-filling estimator, and random estimator, we can replace the $\psi_{m}$ in (I) with ${\hat{p}}_{m}$, ${\hat{n}}_{m}$, and $\frac{Q}{M}$ to obtain ${{\hat{\delta}}_{{\rm{wf}}}}$, ${{\hat{\delta}}_{{\rm{if}}}}$, and ${{\hat{\delta}}_{{\rm{rnd}}}}$, respectively. According to (I), it is obvious that $\hat{\delta}\mathop{=}\limits^{Q\to+\infty}{\cal O}\left({{\sigma^{2}}{M^{2}}{Q^{-1}}}\right)$ holds for the three estimators. This completes the proof.